### v1

# Benchmarking Algorithms for Federated Domain Generalization

Abstract
While prior domain generalization (DG) benchmarks consider train-test dataset heterogeneity, we evaluate Federated DG which introduces federated learning (FL) specific challenges. Additionally, we explore domain-based heterogeneity in clients’ local datasets—a realistic Federated DG scenario. Prior Federated DG evaluations are limited in terms of the number or heterogeneity of clients and dataset diversity. To address this gap, we propose an Federated DG benchmark methodology that enables control of the number and heterogeneity of clients and provides metrics for dataset difficulty. We then apply our methodology to evaluate 13 Federated DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG. Our results suggest that despite some progress, there remain significant performance gaps in Federated DG particularly when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets. Please check our extendable benchmark code here: https://github.com/inouye-lab/FedDG_Benchmark.
1 Introduction
Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on data from an unseen test domain—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization. While most prior DG work focuses on centralized algorithms, another natural context is federated learning (FL) (Konečnỳ et al., 2016), which is a distributed machine learning context that assumes each client or device owns a local dataset. These local datasets could exhibit heterogeneity, which we call client heterogeneity (e.g., class imbalance between clients). Although train-test heterogeneity (in DG) and client heterogeneity (in FL) are independent concepts, both could be naturally defined in terms of domain datasets. For example, suppose a network of hospitals aimed to use FL to train a model to predict a disease from medical images. Because the equipment is different across hospitals, it is natural to assume that each hospital contains data from different domains or environments (or possibly a mixture of domains if it is a large hospital)—this is a case of domain-based client heterogeneity. Yet, the trained model should be robust to changes in equipment within a hospital or to deployment in a new hospital that joins the network—both are cases of domain-based train-test heterogeneity. The interaction between these types of heterogeneity produces new algorithmic and theoretic challenges yet it may also produce new insights and capabilities. Solutions to Federated DG with both types of heterogeneity could increase the robustness and usefulness of FL approaches because the assumptions more naturally align with real-world scenarios rather than assuming the datasets are i.i.d. This could enable training on partial datasets, increase robustness of models to benign spatial and temporal shifts, and reduce the need for retraining.
In the centralized regime, various approaches have been proposed for DG, including feature selection, feature augmentation, etc. Most of these methods are not applicable in the FL regime which poses unique challenges. In the FL regime, client heterogeneity has long been considered a statistical challenge since FedAvg (McMahan et al., 2017), where it experimentally shows that FedAVG effectively mitigate some client heterogeneity. There are many other extensions based on the FedAvg framework tackling the heterogeneity among clients in FL, for example using variance reduction method (Karimireddy et al., 2020). An alternative setup in FL, known as the personalized setting, aims to learn personalized models for different clients to tackle heterogeneity, for example Hanzely et al. (2020). However, none of these works consider model robustness under domain shift between training and testing data. Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains. 2) The evaluations are restricted to the case when the number of clients is equal to the number of domains, which may be an unrealistic assumption (e.g., a hospital that has multiple imaging centers or a device that is used in multiple locations). The case when clients number might be massive are of both theoretical and application interests. 3) None of the works consider the influence of the effect of the number of communication rounds. We provide an overview of the tasks inTable 1, considering both the heterogeneity between training and testing datasets (standard vs. domain generalization) and among clients (domain client heterogeneity). While some studies have addressed the standard supervised learning task, there is a need for a fair evaluation to understand the behavior of domain generalization algorithms in the FL context under those new challenges.
There are several benchmark datasets available for evaluating domain generalization (DG) methods in the centralized setting. These benchmarks, such as DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021), provide multiple datasets that are suitable for assessing the performance of DG algorithms. However, they did not explicitly consider the unique challenges that arise in the federated learning (FL) setting. On the other hand, there are also benchmarks specifically designed for FL. For instance, the LEAF benchmark (Caldas et al., 2018) provides a standardized framework for evaluating FL algorithms. It includes several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario. Another benchmark for FL is PFLBench (Chen et al., 2022), which focuses on evaluating personalized FL methods. PFLBench provides 12 datasets containing various applications. Though these FL-based benchmarks consider statistical heterogeneity, they fail to consider the DG task adequately. Moreover, the level of statistical heterogeneity present in these datasets is insufficient for proper DG evaluation. In summary, DG benchmarks do not consider FL challenges, and FL benchmarks do not consider DG challenges.
Major contributions: We develop a benchmark methodology for evaluating Federated DG with various client heterogeneity contexts and diverse datasets, and we evaluate representative Federated DG approaches with this methodology. 1) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 3). 2) We propose a standardized definition of domain-based client heterogeneity that is unique to the FL context and interpolates between domain homogeneity and domain separation (see Fig. 1) while controlling class imbalance. In particular, we develop a novel method to split any dataset with domain labels across any number of clients (see subsection 3.2). 3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG. Our results indicate that there still exist significant gaps and open research directions in Federated DG. 4) We will release an extendable open-source library for evaluating Federated DG methods.
Notation
[…]
2 Approaches to Federated Domain Generalization
We first briefly review the DG problem, extend to the Federated DG problem, and explain domain-based client heterogeneity. Then, we discuss three categories of Federated DG approaches.
2.1 Problem Background and Setup
Domain generalization (train-test heterogeneity)
Unlike standard ML which assumes the train and test are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of the test domain distributions when only samples from the train domain distributions are given. Formally, given a set of train domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
$\min_{\theta}{\textstyle\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{d\in\mathcal{D}_{\text{test}}}}\mathcal{L}(\theta;p_{d})\quad\text{or}\quad\min_{\theta}\max_{d\in\mathcal{D}_{\text{test}}}\mathcal{L}(\theta;p_{d})\,,$ (1)
where $\mathcal{L}(\theta;p_{d})=\mathbb{E}_{(x,y)\sim p_{d}}[\ell(x,y;\theta)]$ where $\ell$ is a per-sample loss function such as squared or cross-entropy loss. The key challenge in DG is that the train and test domain distributions are disjoint, i.e., $\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$, and thus, the method must be able to generalize beyond the train domain distributions to perform well on the test domain distributions. The naïve approach is to simply ignore the domains and perform empirical risk minimization (ERM) on all the training data—which is actually a challenging baseline to outperform in practice (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021).
Federated DG
Federated DG adds a layer of complexity because now the domain distribution samples are not centrally located. Instead, each FL client can update their local model based only on their local client distribution $p_{c}$ and then pass their local parameters $\theta_{c}$ to a central server, which will aggregate the local models and broadcast the model to all clients. The FL problem can be abstracted as follows:
$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Optimize}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\quad\text{and}\quad\underbrace{\theta_{\text{global}}=\texttt{Aggregate}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
where the client distributions may be homogeneous (i.e., $\forall(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Optimize minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$, and Aggregate aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
Domain-based client heterogeneity
While client heterogeneity (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client. At one extreme, FL with i.i.d. data would be equivalent to the mixture proportions being the same across all clients, i.e., $\forall c,c^{\prime},d,w_{c,d}=w_{c^{\prime},d}$, which we call the homogeneous setting. On the other extreme, if the number of clients and number of domains are equal ($D=C$), the domain weights could be disjoint between clients such that each client “owns” one or more domains, i.e., for any domain $d\in[D],$ $w_{c,d}>0\Leftrightarrow w_{c^{\prime},d}=0,\forall c^{\prime}\neq c$, which we call domain separation (in subsection 3.2, we extend the domain separation to the case $D\neq C$). Finally, we use $\lambda$ to denote an interpolation parameter between these two extremes: homogeneous case $\lambda=1$ and domain separation case $\lambda=0$. We give details of an explicit algorithm to use in practice for splitting datasets for evaluation in subsection 3.2 which covers two extremes and interpolation cases $\lambda\in(0,1)$. Table 1 summarizes the train-test heterogeneity in DG and the domain-based client heterogeneity from the FL context, where we focus on Federated DG.
Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the train and test datasets (columns). This paper focuses on the DG setting (right column).
Between train and test datasets (train-test heterogeneity)
Standard supervised learning Domain generalization (our focus)
Among clients Homogeneous ($\lambda=1$) FL with homogeneous clients Federated DG with homogeneous clients
Heterogeneous ($10<\lambda<$ 1) FL with heterogeneous clients Federated DG with heterogeneous clients
Domain separation ($\lambda=0$) FL with domain separation Federated DG with domain separation
Overview of Federated DG methods
In this benchmark study, we explore three categories of Federated DG methods: DG methods originally designed for the centralized setting, FL methods specifically tailored to handle client heterogeneity, and methods specifically designed for Federated DG. To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories and compare to vanilla FedAvg (McMahan et al., 2017) with ERM loss, where any heterogeneity is simply ignored. These methods are selected based on their prominence and potential effectiveness for Federated DG. To ensure a diverse range of evaluation scenarios, we conduct experiments on various datasets and under various heterogeneity settings. Our results provide insights into their relative performance and suitability for different Federated DG contexts. This benchmark study aims to offer a broad overview of the current popular methods, enabling researchers and practitioners to make informed decisions when tackling Federated DG.
2.2 Centralized DG methods adapted to the FL setting
The first natural choice is directly migrating the DG methods from the centralized regime. To adapt those methods, we simply run the centralized DG method at each client locally with their own local dataset (see subsection 3.2 for how the local datasets are created), and then compute an average of model parameters at each communication round (see next paragraph). This approach is straightforward for the homogeneous ($\lambda=1$) and heterogeneous ($\lambda=0.1$) settings where each client has data from all training domains—albeit quite imbalanced for $\lambda=0.1$. This can be seen as biased updates at each client based on biased local data. In the domain separation case $\lambda=0$, (i.e., if all clients only have one primary domain, i.e., $\forall k,|P_{k}|=1$), this simple approach cannot be applied because centralized DG methods require data from at least two domains. In fact, these centralized DG methods degenerate to vanilla FedAvg if there is only one domain per client. Extending these methods to the case where all clients only have one domain without violating the FL constraints is an interesting direction for future work.
A predominant and effective centralized DG approach is through representation learning, including domain-invariant representation learning by learning domain-invariant features via either kernel methods (Muandet et al., 2013; Ghifary et al., 2016) or invariant risk minimization (Arjovsky et al., 2019; Krueger et al., 2020), and domain adversarial neural networks Sun and Saenko (2016), and Li et al. (2018b) which explicitly tries to align the feature representation distribution. Besides invariant representation, there are other methods the general learning strategy to promote the generalization ability gradient operation Shi et al. (2021); Rame et al. (2022), which tries to learn generalized representations by directly operating on gradients. Other approaches include distributionally robust optimization Sagawa et al. (2019), which learns the worst-case distribution scenario of training domains; and meta-learning Finn et al. (2017); Li et al. (2018a), which is based on the learning-to-learn mechanism to learn general knowledge by constructing meta-learning tasks to simulate domain shift. We selected IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), MMD (Li et al., 2018b), Mixup (Zhang et al., 2017), DeepCoral (Sun and Saenko, 2016), and GroupDRO (Sagawa et al., 2019) from this category for their representative mechanisms.
2.3 FL methods tackling client heterogeneity
Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task). The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al. (2018). Several subsequent works have analyzed client heterogeneity by assuming bounded gradients (Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019) or bounded gradient dissimilarity (Li et al., 2020), and additionally assuming bounded Hessian dissimilarity (Karimireddy et al., 2020; Khaled et al., 2020; Liang et al., 2019). From this category, we selected FedProx (Li et al., 2020), which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model. Scaffold (Karimireddy et al., 2020) utilizes variance reduction to account for client heterogeneity.
2.4 FL methods designed for Federated DG
Limited research has focused explicitly on solving the Federated DG by design. FedDG Liu et al. (2021) introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol. Another approach, FedADG (Zhang et al., 2021), utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator. FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach. However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging. A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) (Tenison et al., 2022) aims to enhance generalization across clients and the global model. FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components. FedSR (Nguyen et al., 2022) proposes a simple algorithm that uses two locally-computable regularizers for domain generalization. Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
3 Benchmark Methodology
In this study, we aimed to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup. We evaluated a total of $13$ methods, encompassing three different types of approaches.
3.1 Four evaluation dimensions
(1) Domain-based client heterogeneity and (2) number of clients
Previous studies on Federated DG have often focused on domain separation client heterogeneity where the number of clients equals the number of training domains, i.e., $C=D$. However, this excludes evaluation of the homogeneous and partially heterogeneous settings and restricts the number of clients to the number of training domains $D$. In particular, many pseudo-realistic datasets, such as those in DomainBed (Gulrajani and Lopez-Paz, 2020), which consist training domains $D\leq 5$, which limits the evaluation of methods under scenarios with a large number of clients. Conversely, when using realistic domain datasets that have many domains, such as those in WILDS (Koh et al., 2021), most current methods perform poorly under this extremely challenging setting. By introducing the domain split method discussed in subsection 3.2, we can explore various levels of client heterogeneity and relax the assumption that $C=D$ so that we can leverage both pseudo-realistic and realistic datasets and evaluate methods at an appropriate difficulty level.
(3) Dataset difficulty and (4) dataset type
While most Federated DG work focuses on standard image-based datasets, we evaluate methods across a broad range of dataset difficulty (ranging from easy pseudo-realistic datasets to very challenging realistic datasets) and two types of datasets ($3$ image datasets and $2$ text datasets). This ensures that we can fully understand the performance of each method across a wide range of scenarios. Further details can be found in Section subsection 3.3.
3.2 Domain-based client heterogeneity by splitting DG datasets
For evaluation dimensions (1) and (2) above, we need to control over the amount of client heterogeneity, ranging from homogeneous to domain separation, and we need to control over the number of clients, which may be smaller or larger than than the number of domains (i.e., $D\neq C$). We propose a way to split any DG dataset into any number of clients that allows explicit control of the amount of heterogeneity through the $\lambda$ hyperparameter which attempts to balance the number of samples per client. An illustration of our domain splitting procedure can be found in the following Fig. 1.
Figure 1: Distinct color refers to distinct domain data, and $\lambda$ is the domain balancing parameter. (a): Train-test domain heterogeneity. (b): Domain split when $C\leq D$ and domain split when $C>D.$ (c): domain split when domain number exceeds client number, i.e., $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and separation ($\lambda=0$). See Algorithm 1 for domain split when $C>D.$
Algorithm 1 DomainSplit Algorithm
Input Number of clients $C$, heterogeneity parameter $\lambda\in[0,1]$, and samples from each domain $(\mathcal{S}_{1},\cdots,\mathcal{S}_{D})$, where $n_{1},\dots,n_{D}$ denote the number of samples per domain and w.l.o.g. are assumed to be in descending order, i.e., $n_{1}\geq n_{2}\geq\dots\geq n_{D}$.
  // Divide domain indices across clients
  if $C\leq D$ then
     $\forall c,\mathcal{D}_{c}\leftarrow\emptyset$
     for $d=1,2,\dots,D$ do
        // Find client with fewest samples
        $c^{*}\in{\textstyle\operatorname*{arg\,min\,}\limits_{c\in[C]}\,\sum_{d^{\prime}\in\mathcal{D}_{c}}}n_{d^{\prime}}$
        $\mathcal{D}_{c^{*}}\leftarrow\mathcal{D}_{c^{*}}\cup\{d\}$
     end for
  else if $C>D$ then
     $\forall c\in\{1,2,\dots,D\},\,\mathcal{D}_{c}\leftarrow\{c\}$
     for $c=D+1,\dots,C$ do
        // Find on average largest domain to split
        $d^{*}\in\operatorname*{arg\,max\,}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}$
        $\mathcal{D}_{c}\leftarrow\{d^{*}\}$
     end for
  end if
  // Split samples across the clients
  $\forall c,\mathcal{S}_{c}\leftarrow\emptyset$
  for $c\in[C],d\in[D]$ do
     $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1-\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]\cdot n_{d}}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}$
     $\mathcal{S}^{\prime}\leftarrow\texttt{SampleWOReplacement}(\mathcal{S}_{d},n_{d,c})$
     $\mathcal{S}_{d}\leftarrow\mathcal{S}_{d}\setminus S^{\prime}$// Remove from domain
     $\mathcal{S}_{c}\leftarrow\mathcal{S}_{c}\cup S^{\prime}$// Append to client dataset
  end for
Output $(\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{C})$
In Algorithm 1, we provide a concrete algorithm for splitting samples from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via $\lambda$, ranging from homogeneous clients to domain separation. Our algorithm has two main steps. In Step 1, we assign “primary” domain indices $\mathcal{D}_{c}\subseteq[D]$ to each client $c\in[C]$ depending on $C$ and $D$. If $C\leq D$, the domains are sorted in descending order according to number of sample size and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples (denoted by $\sum_{d^{\prime}\in\mathcal{D}_{c^{*}}}n_{d^{\prime}}$). In this case, the algorithm ensures that no client shares domains with the others but otherwise attempts to balance the total number of training samples between clients. If $C>D$, we first assign the domains one by one to the first $D$ clients. Then, starting from client $c=D+1$, we iteratively assign the current on average largest domain $d^{*}$ to $c,$ accounting for the fact that the domain may already shared by different clients—notationally, on average is represented through dividing by $\sum_{c^{\prime}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]$. In this case, some clients may share one domain, but no client holds two domains simultaneously while otherwise attempting to balance the number of samples across clients as much as possible.
In Step 2, we define the sample counts for each client and domain, denoted $n_{c,d}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$ $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1\!\!-\!\!\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}n_{d}\,,$ where rounding to integers is carefully handled when not perfectly divisible and where $\mathbbm{1}[\cdot]$ is the indicator function. This is simply a convex combination between homogeneous clients ($\lambda=1$) and domain separation ($\lambda=0$). Given the number of samples per client per domain, we simply sample without replacement from the corresponding domain datasets and build up the client datasets.
3.3 Dataset Selection and Dataset Difficulty Metrics
To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities. We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context. For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance with and without samples from the test domain (i.e., the former is able to “cheat” by seeing part of test domain samples during training). For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples. These dataset difficulty metrics can be formalized as follows:
$R_{\text{DG}}$ $\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})}$
$R_{\text{FL}}(\lambda)$ $\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the train domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ split respectively of $\mathcal{S}_{\text{DG-test}}$. For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the train domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from DG effect. We choose five datasets in our benchmark: FEMNIST from Caldas et al. (2018), PACS from Gulrajani and Lopez-Paz (2020), and IWildCam, CivilComments and Py150 from Koh et al. (2021). We summarize the statistics and difficulty metrics in Table 2 and provide the rationale of selecting these datasets in the Appendix.
Table 2: Summary of selected datasets as well as their difficulty metric, where $n$ is the number of samples, $D$ is the number of domains, $C$ is the number of clients used.
Dataset Statistics Difficulty Metric
$n$ $D$ $C$ $R_{\text{DG}}$ $R_{\text{FL}}(\lambda)$
$\lambda=1$ $\lambda=0.1$ $\lambda=0$
FEMNIST $737036$ $2586$ $100$ $1.000$ $0.980$ $0.981$ $0.981$
PACS $9991$ $2$ $100$ $0.960$ $1.000$ $1.000$ $1.000$
IWildCam $203029$ $243$ $243$ $0.449$ $0.682$ $0.525$ $0.525$
CivilComments $448000$ $16$ $100$ $0.984$ $0.618$ $0.533$ $0.532$
Py150 $150000$ $8421$ $100$ $0.969$ $0.999$ $0.998$ $0.998$
4 Benchmark Experimental Results
In this section, we report the performance of $13$ representative methods from three lines of research on $5$ different datasets, where the FEMINIST results are provided in the appendix given the DG task simplicity (i.e., $R_{\text{DG}}\approx 1$). For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison. We then select the model based on the held-out-domain validation set. After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3 and Table 4. See Appendix for detailed hyperparameters choices. We also include the results according to in-domain validation early-stopping in the Appendix but the trends are similar. We make the following remarks on the main results from Table 3 and Table 4.
Remark 4.1.
FedAvg with an ERM objective is a strong baseline. Simple FedAvg with an ERM objective is a strong baseline that is challenging to beat across datasets (except for CivilComments), similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021). We recommend always including FedAvg as a baseline in all future evaluations.
Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.8379$ $0.8766$ $0.9144$ $\mathit{0.9377}$ $0.2727$ $0.1762$ $0.1241$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $\mathit{0.3134}$ $0.0718$ $0.0924$
DeepCoral $\bf 0.9216$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\bf 0.9395$ $\bf 0.9437$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8826$ $0.8407$ $0.8892$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8796$ $0.9239$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.9234}$ $\mathit{0.9275}$ $\mathbf{0.9521}$ $0.1445$ $0.1191$ $0.0503$
FedADG - $0.1892$ $0.0592$ $0.0598$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1246$ $0.1263$ $0.1257$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8446$ $0.0106$ $0.0957$ $0.0106$
Table 4: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $\mathit{0.6405}$
DG Adapted IRM $0.6408$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $\mathbf{0.6713}$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $\mathit{0.6520}$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $0.6565$
DeepCoral $0.5847$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $0.6383$ $0.6028$ $0.4954$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - 0.3603 0.3389 0.3188 - 0.0915 0.0860 0.0587
FedGMA - 0.3341 0.3157 0.3195 - 0.6628 0.3737 0.3576
Remark 4.2.
Most centralized DG methods degrade in the FL setting. For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3. Further, degradation can be seen in PACS when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
Remark 4.3.
FL methods tackling client heterogeneity perform surprisingly well on the Federated DG task. FedProx and Scaffold, which were designed for client heterogeneity but not specifically for the DG task, perform quite well in the Federated DG setting and even perform the best for IWildCam and Py150. Thus, we suggest including these methods in future Federated DG evaluations and suggest that future works could focus on combining the strengths of these methods and DG methods.
Remark 4.4.
The performance of real-world data significantly degrades as $\lambda$ decreases. This can be seen from IWildCam and Py150. While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG. We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
Remark 4.5.
A diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG. This conclusion is inspired from the following two observations. 1) When comparing FL methods, we notice opposite trends between PACS and all other datasets (IWildCam, CivilComments, Py150) when increasing client heterogeneity based on $\lambda$. For PACS, the FL methods (except for FedADG and FedSR) surprisingly seem to improve when increasing client heterogeneity (i.e., $\lambda=0$). However, for other datasets, the expected trend exists such that increasing client heterogeneity produces worse performance. From this, we recommend that using PACS alone might be a misleading or at least incomplete evaluation of Federated DG methods. 2) Centralized DG methods adapted to FL perform the best on CivilComments. This unexpected result suggests that DG methods in the centralized regime may be able to better accommodate subpopulation shift, which is the special kind of shift that is exhibited in CivilComments. Both of these observations emphasize that a diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG tasks.
Additional DG challenges from FL.
For further understanding, we explore some additional questions on the PACS dataset because it is the most common and computationally feasible to train many different models. Specifically, we explore how the number of clients, amount of communication (i.e., the number of server aggregations) in the federated setup, and client heterogeneity affects the performance of various methods. The figures and detailed analysis are provided in the Appendix but we highlight two remarks here.
Remark 4.6.
The number of clients $C$ strongly affects on the overall performance. The DG performance drops from $90\%$ to $50\%$ or even 10% when varying $C$ from $1$ to $200$. We strongly recommend future Federated DG evaluations to consider larger number of clients such as $50$ to $100$ rather than only a very small number of clients. It indicates that there exist significant unresolved challenges in the domain of federated domain generalization when dealing with a large number of clients, which have not been adequately explored. In particular, FedADG and FedSR seem to be sensitive to the number of clients because they perform poorly when $C\geq 10$, while in the original papers they were only evaluated on a few clients $C=3$ (and we reproduce the good results with such $C$ in the Appendix).
Remark 4.7.
The number of communications does not monotonically affect DG performance. We observe an interesting implicit regularization phenomena in the FL context: different methods achieve their best performance when the communication rounds is relatively low. In particular, for PACS, this communication round is $10$ (while fixing the total amount of computations). This is surprising as in contrast for in-domain task, FL suggests more communications leads to better performance (Stich, 2018, Theorem 5.1). Further theoretical investigation of the dependence of DG accuracy on the communications rounds, and its possible relation to implicit regularization via early stopping is an interesting area for future work.
5 Conclusion and Discussion
We first build a systematic benchmark methodology for evaluating Federated DG that includes novel methodologies for splitting the data across clients and evaluation of dataset difficulty. We then evaluate $13$ representative methods from three relevant lines of research. Our evaluation shows that Federated DG is still unsolved and significant gaps remain between centralized DG performance and Federated DG performance. Therefore, Federated DG is ripe for additional research. Based on our evaluation and observation, here are some recommendations and suggestions for future work in Federated DG.
Recommendations for future evaluations of Federated DG.
  1. 1.
    Stronger Baselines - FedAvg should always be included because it is a strong baseline (Remark 4.1). FL methods designed for client heterogeneity (though not necessarily DG) should also be included given their strong performance (Remark 4.3).
  2. 2.
    Realistic Datasets - Federated DG methods should be evaluated on more realistic datasets. In particular, FEMNIST and PACS behave quite differently with respect to client heterogeneity than the more realistic WILDS datasets (Remark 4.4). Thus, we recommend including both IWildCam and Py150 in future evaluations of Federated DG. CivilComments may be useful for evaluating realistic subpopulation shift (Remark 4.5).
  3. 3.
    Large Number of Heterogeneous Clients - Evaluations should include scenarios with a large number of clients with varying degrees of heterogeneity. A large number of clients poses unique challenges for most methods (Remark 4.6). Additionally, the heterogeneity of clients is more realistic and plays a significant role in evaluation (Remark 4.4).
Suggestions for future work in Federated DG.
  1. 1.
    Handle Domain Separation Case - The domain separation scenario ($\lambda=0$) limits the exchange of information between domains, i.e., a client may only have data from a single domain. Centralized approaches cannot be adapted to this setting and current methods still struggle under this realistic client heterogeneity setting.
  2. 2.
    Increase Convergence Rate - In most benchmark experiments (except for PACS, which is an easy dataset), we have observed slow convergence, particularly on challenging real-world datasets such as IWildCam and Py150. Moreover, we have noticed even slower convergence when client heterogeneity is high (i.e., when $\lambda$ is small). In addition, we have found that federated methods designed to improve convergence actually performed quite well even though they were not designed for the DG task (Remark 4.3). Therefore, it is crucial for future methods to improve the convergence rate of Federated DG approaches.
  3. 3.
    Investigate the Effect of Communication Frequency - Because increased communication frequency may actually hurt DG performance (Remark 4.7), further investigation into the regularization effect of early stopping and infrequent communications in FedAvg may yield important insights.
  4. 4.
    Understand the Effect of the Number of Clients - Given that the performance of some methods degrade very quickly when increasing the number of clients beyond a few (Remark 4.6), the community could benefit from an improved theoretic and empirical understanding of how the number of clients affects Federated DG in terms of convergence, computational requirements, sample complexity, and DG performance.
We hope this work provides a better foundation for future work in Federated DG and accelerates research progress.
Acknowledgements
This work was supported by Army Research Lab under Contract No. W911NF-2020-221. R.B. and D.I. also acknowledge support from NSF (IIS-2212097) and ONR (N00014-23-C-1016). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
Appendix
[…]
Reproducibility Statement
Code for reproduce the results is available at the following link:
https://github.com/inouye-lab/FedDG_Benchmark.
We include detailed documentation in using our code to reproduce the results throughout the paper. We also provide documentation in adding new algorithm’s DG evaluation in the FL context.
Appendix A Datasets and Difficulty Metric
A.1 Dataset Introduction
In this section, we introduce the datasets we used in our experiments, and the split method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
FEMNIST
It is an FL prototyping image dataset of handwritten digits and characters each users created as a natural domains, widely used for evaluation for client heterogeneity in FL. Though it contain many training domains, it lacks significant distribution shifts across domains ($R_{\text{DG}}=1$), and considered as easy compared to other datasets.
PACS
It is an image dataset for domain generalization. It consists of four domains, namely Photo (1,670 images), Art Painting (2,048 images), Cartoon (2,344 images), and Sketch (3,929 images). This task requires learning the classification task on a set of objects by learning on totally different renditions. $R_{\text{DG}}=0.960$ makes it a easy dataset as domain generalization in our setting. Notice that choosing different domain as test domain might give us different $R_{\text{DG}}$.
IWildCam
It is a real-world image classification dataset based on wild animal camera traps around the world, where each camera represents a domain. It contains 243 training domains, 32 validation and 48 test domains. Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference Koh et al. [2021]. The $R_{\text{DG}}=0.449$ makes it a very challenging dataset for domain generalization.
CivilComments
It is a real-world binary classification text-based dataset formed from the comments from different demographic groups of people, containing $8$ demographic group. The goal is to judge whether the text is malicious or not.
Py150
It is a real-world code-completion dataset which is challenging given massive domains $8421,$ where the domain is formed according to the repository author. The goal is to predict the next token given the context of previous tokens. We evaluate models by the accuracy on the class and method tokens.
A.2 Dataset Split Setup
For each dataset, we first split the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset. For FEMNIST and PACS dataset, we use Cartoon and Sketch as training domain, Art-painting as held-out domain, Painting as test domain. For training domain, we split $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively. For IWildCam, CivilComments and Py150, we directly apply the Wilds official splits.
Appendix B Benchmark Experimental Setting
B.1 Model Structure
In this benchmark, for image-based datasets, we use ResNet-50 He et al. [2016] dataset. For CivilComments and Py150 dataset, we choose DistilBERT Sanh et al. [2020] and CodeGPT Radford et al. [2019] respectively as recommended by Wilds.
B.2 Model Selection
We conduct held-out domain model selection with $4$ runs for each methods. The oracle model selection evaluates the model based on the performance on the held-out validation domain. The results are reported based on the best run.
B.3 Early Stopping
We conduct early stopping using the held-out validation dataset in our evaluation. For each dataset and method, We first run certain communication rounds, then we select the model parameters which achieves the best performance on the validation set. We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix D.
B.4 Hyperparameters
In this section, we present the hyperparameters selected for the evaluation. We opted to communicate per epoch for all experiments. For PACS, we run totally $50$ communication rounds. For IWildCam, we run $50$ communication rounds. For CivilComments, we run $10$ communication rounds. For Py150, we run $10$ communication rounds, and for FEMNIST, we run $40$ communication rounds. Please refer to the Table 5 to review other hyperparameters.
Table 5: Hyperparameters used in the evaluation. c represents centralized learning regime, and f represents the federated learning regime
PACS IWildCam CivilComments Py150 FEMNIST
c f c f c f c f c f
Fed Avg Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
IRM Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty weight $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$
Penalty anneal iters $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$
Fish Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
meta learning rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
Mixup Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\alpha$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$
MMD Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Deep Coral Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Penalty Weight $1$ $1$ $10$ $10$ $1$ $1$ $1$ $1$ $1$ $1$
Group DRO Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\eta$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$
Fed Prox Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\mu$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
Scaf fold Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
FedDG Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
FedADG Classifier optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGDs
Classifier learning rate $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$
Gen optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Gen learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
Disc optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
Disc learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
$\alpha$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$
FedSR Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
$\ell_{2}$ regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.001$ $0.001$ $0.01$ $0.01$
cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
FedGMA Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
Server step size $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
Mask threshold $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
Appendix C Additional FL-specific challenges for domain generalization
As mentioned in Section 4, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $K$ increases on PACS. We fix the communication rounds $50$ and the local number of epoch is $1$ (synchronizing the models every epoch). Fig. 2 plots the held-out DG test accuracy versus number of clients for different levels of data heterogeneity. The following comments are in order: given communication budget, 1) current domain generalization methods all degrade a lot after $C\geq 10$, while the performance ERM and FedDG maintain relatively unchanged as the clients number increases given communication budget. FedADG and FedSR are are sensitive to the clients number, and they both fail after $C\geq 20.$ 2) Even in the simplest homogeneous setting $\lambda=1,$ where each local client has i.i.d training data, current domain generalization methods IRM, FISH, Mixup, MMD, Coral, GroupDRO work poorly in the existence of large clients number, this means new methods are needed for DG in FL context when data are stored among massive number of clients.
Figure 2: PACS: Held-out DG test accuracy versus number of clients.
ii) Communication constraint: To show the effect of communication rounds on convergence, we plot the test accuracy versus communication rounds in Appendix Fig. 3. We fix the number of clients $C=100$ on PACS and decreases rounds of communication (together with increasing local epochs). That is, if the regime restricts the communication budget, then we increase its local computation $E$ to have the same the total computations. Therefore, the influence of communication on the performance is fair between algorithms because the total data pass is fixed. We observe that the total number of communications does not monotonically affect the DG performance. With decreasing number of total communication, most methods’ performance first increase then decrease. This might be an interesting implicit regularization phenomena in the FL context. Without discussing DG task, usually frequent communications lead to faster convergence. The relationship between DG performance and communications requires further exploration.
Figure 3: PACS: Held-out DG test accuracy vs. varying communications (resp. varying echoes ).
Appendix D Supplementary Results
In the main paper, we provide experiments results using held-out validation early stopping. Here we report the results using in-domain validation set. We also report the convergence curve for each methods on each dataset for reference. We observe that under most of cases, the held-out validation gives us a better model. Thus, we recommend using held-out validation set to perform early stopping.
More results on PACS and IWildCam dataset.
Here we both report the convergence curve using the DG accuracy of each method on PACS Fig. 4 and IWildCam Fig. 5 and the DG accuracy on PACS and IWildCam Table 6 using the in-domain validation. From the PACS convergence curve Fig. 4, we observe that with lower $\lambda$, the model is harder to converge. FedAvg ang FedDG converge faster than all other methods while FedADG and FedSR does not converge. From IWildCam convergence curve Fig. 5, we observe that all methods are struggling to converge. This is due to the main challenge first come from the $R_{\text{FL}}$ here. We thus observe that Federated methods dealing with heterogeneity clients achieves best performance. Especially when $\lambda$ is low. It is worth noticing that this does not mean DG in FL is solved, none of the methods even achieve the centralized ERM performance.
Figure 4: Convergence curve on PACS; total clients and training domains $(C,D)=(100,2);$ increasing domain heterogeneity from left to right: $\lambda=(1,0.1,0).$
Figure 5: Accuracy versus communication rounds for IWildCam; Total clients number $C=243$; increasing heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
Table 6: Test accuracy on PACS and IWildCam dataset with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
PACS ($D=2$) IWildCam ($D=243$)
$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.8389$ $0.8766$ $\mathit{0.9144}$ $\mathbf{0.8958}$ $\mathit{0.2727}$ $\mathbf{0.1762}$ $\mathbf{0.1241}$ $0.0707$
DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $0.3134$ $0.0718$ $0.0924$
DeepCoral $\mathbf{0.9216}$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
GroupDRO $0.9060$ $\mathbf{0.9395}$ $\mathbf{0.9437}$ $0.2531$ $0.0530$ $0.0908$
FL FedProx - $0.8325$ $0.8407$ $0.8568$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
Scaffold - $0.8215$ $0.8643$ $0.8787$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
FDG FedDG - $\mathit{0.8868}$ $0.8880$ $\mathit{0.8952}$ $0.1445$ $0.1023$ $0.0503$
FedADG - $0.0915$ $0.0732$ $0.0592$ $0.0049$ $0.0049$ $0.0049$
FedSR - $0.1180$ $0.1341$ $0.1138$ $0.0056$ $0.0056$ $0.0056$
FedGMA - $0.8820$ $0.8467$ $0.8180$ $0.0271$ $0.0259$ $0.0081$
More results on CivilComments and Py150 dataset.
Here we report the DG accuracy on CivilComments and Py150 Table 6 using the in-domain validation. We could observe that most of the methods perform well on CivilComments and Py150. This could be attributed to the use of pretrained models. Additionally, the utilization of pretrained models may explain the relatively high value of $R_{\text{DG}}$, as NLP pretrained models have already been exposed to various domains.
Table 7: Test accuracy on CivilComments and Py150 datasets with in-domain validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column. FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
CivilComments ($D=16$) Py150 ($D=5477$)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $0.6405$
DG Adapted IRM $\mathit{0.6213}$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
Fish $0.6038$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
MMD $0.5981$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $\mathit{0.6565}$
DeepCoral $0.5827$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
GroupDRO $\mathbf{0.6383}$ $0.6028$ $0.4804$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
FDG FedSR - $0.3603$ $0.3389$ $0.3188$ $0.0915$ $0.0860$ $0.0587$
FedGMA - $0.3341$ $0.3157$ $0.3195$ $0.6628$ $0.3737$ $0.3576$
Results on FEMNIST dataset.
As mentioned in the main paper, we include the FEMNIST dataset here as reference. We could observe from Table 8 that $\lambda$ does not influence the final DG accuracy, and either in-domain validation or held-out-domain validation does not affect the final DG accuracy. This indicates the lacking of the statistical heterogeneity across different domains. we observe that changing of $\lambda$ does not significantly affect the convergence. Most of which does not converge to the centralized counterpart’s performance, this is due to the challenge coming from large number of clients where $R_{\text{FL}}=0.980<1$.
Figure 6: Accuracy versus communication rounds for FEMNIST; total clients and training domains $(K,M)=(100,2586);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
Table 8: Test accuracy on FEMNIST dataset with in-domain validation and held-out validation where FedAvg-ERM is the simple baseline (B). “-” means the method is not applicable in that context. Bold is for best and italics is for second best in each column.
FEMNIST (In-Domain) FEMNIST (Held-Out-Domain)
$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
B
FedAvg-ERM $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$ $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$
DG Adapted IRM $0.8442$ $0.8303$ $0.8389$ $0.8328$ $0.8442$ $0.8303$ $0.8389$ $0.8328$
Fish $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$ $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$
Mixup $0.8337$ $0.8275$ $0.8168$ $0.8161$ $0.8337$ $0.8275$ $0.8168$ $0.8161$
MMD $0.8442$ $0.8430$ $0.8323$ $0.8293$ $0.8442$ $0.8430$ $0.8323$ $0.8293$
DeepCoral $0.8457$ $0.8403$ $0.8328$ $0.8513$ $0.8457$ $0.8403$ $0.8328$ $0.8513$
GroupDRO $0.8413$ $0.8447$ $0.8139$ $0.8052$ $0.8413$ $0.8447$ $0.8139$ $0.8052$
FL FedProx - $0.8448$ $0.8440$ $0.8425$ $0.8448$ $0.8440$ $0.8425$
Scaffold - $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$ $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$
FDG FedDG - $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$ $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$
FedADG - $0.0549$ $0.0557$ $0.0548$ $0.0549$ $0.0557$ $0.0548$
FedSR - $0.8320$ $0.8316$ $0.8306$ $0.8320$ $0.8316$ $0.8306$
FedGMA - $\mathit{0.8486}$ $0.8395$ $0.8371$ $0.8486$ $0.8395$ $0.8371$
Appendix E Gap Table
We list the gap table in Table 9 for summarizing the current DG algorithms performance gap w.r.t FedAvg-ERM in the FL context, in particular, positive means it outperforms FedAvg-ERM, negative means it is worse than FedAvg-ERM. It can be seen that in the on the simple dataset, the best DG migrated from centralized setting is better than FedAvg-ERM. In the domain separation case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting. However, they fail in harder datasets. In the hardest setting, currently the Federated methods dealing with data heterogeneity performs the best. It is worth noting that while federated learning methods that address client heterogeneity perform better than other methods, they still fall short of achieving centralized empirical risk minimization (ERM). This highlights the need for future research and development of DG methods in the FL regime.
Table 9: Gap Table: Current progress in solving Federated DG with respect to FedAvg-ERM.
Domain Separation
Free of Data Leakage
PACS
IWildCam
CivilComments
Py150
Adapted centralized methods $+0.0629$ $-0.0981$ $+0.0421$ $-0.0062$
Federated methods $+0.0060$ $+0.0458$ $+0.0421$ $-0.0062$
FDG FedDG $+0.0468$ $-0.0775$
FedADG $-0.6874$ $-0.2171$
FedSR $-0.7520$ $-0.2164$ $-0.0326$ $-0.5912$
FedGMA $+0.0054$ $-0.1263$ $-0.0588$ $-0.0199$
Appendix F Training Time, Communication Rounds and local computation
In this section, we provide training time per communication in terms of the wall clock training time. Notice that for a fixed dataset, most of algorithms have similar training time comparing to FedAvg-ERM, where FedDG and FedADG are significantly more expensive.
Table 10: Wall-clock Training time per communication (unit: s).
PACS IWildCam CivilComments Py150 FEMNIST
$C=100$ $C=243$ $C=100$ $C=100$ $C=100$
B
FedAvg-ERM $143$ $6301$ $3958$ $6566$ $262$
DG Adapted IRM $147$ $6454$ $4085$ $7089$ $297$
Fish $148$ $7072$ $5483$ $7770$ $324$
Mixup $144$ $6294$ $264$
MMD $144$ $6663$ $4024$ $7603$ $287$
DeepCoral $144$ $6597$ $3901$ $7212$ $287$
GroupDRO $145$ $9311$ $4690$ $8121$ $307$
FL FedProx $169$ $6921$ $4502$ $6513$ $288$
Scaffold $167$ $6876$ $4421$ $6389$ $281$
FDG FedDG $352$ $32172$ $989$
FedADG $181$ $11094$ $907$
FedSR $151$ $7136$ $4403$ $8020$ $280$
FedGMA $143$ $6795$ $4525$ $6545$ $261$

### v2

@@
 Abstract
-While prior domain generalization (DG) benchmarks consider train-test dataset heterogeneity, we evaluate Federated DG which introduces federated learning (FL) specific challenges.
-Additionally, we explore domain-based heterogeneity in clients’ local datasets—a realistic Federated DG scenario.
-Prior Federated DG evaluations are limited in terms of the number or heterogeneity of clients and dataset diversity.
-To address this gap, we propose an Federated DG benchmark methodology that enables control of the number and heterogeneity of clients and provides metrics for dataset difficulty.
-We then apply our methodology to evaluate 13 Federated DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG.
-Our results suggest that despite some progress, there remain significant performance gaps in Federated DG particularly when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets.
-Please check our extendable benchmark code here: https://github.com/inouye-lab/FedDG_Benchmark.
+While prior federated learning (FL) methods mainly consider client heterogeneity, we focus on the Federated Domain Generalization (DG) task, which introduces train-test heterogeneity in the FL context.
+Existing evaluations in this field are limited in terms of the scale of the clients and dataset diversity.
+Thus, we propose a Federated DG benchmark that aim to test the limits of current methods with high client heterogeneity, large numbers of clients, and diverse datasets.
+Towards this objective, we introduce a novel data partition method that allows us to distribute any domain dataset among few or many clients while controlling client heterogeneity.
+We then introduce and apply our methodology to evaluate $14$ DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG on $7$ datasets.
+Our results suggest that, despite some progress, significant performance gaps remain in Federated DG, especially when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets.
+Furthermore, our extendable benchmark code will be publicly released to aid in benchmarking future Federated DG approaches.
 1 Introduction
-Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on data from an unseen test domain—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization.
+Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on test data coming from a different distribution from training distribution—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization.
 While most prior DG work focuses on centralized algorithms, another natural context is federated learning (FL) (Konečnỳ et al., 2016), which is a distributed machine learning context that assumes each client or device owns a local dataset.
@@
 (2020).
+There are also unsupervised FL methods tackling domain translation instead of classification (Zhou et al., 2022; Wang et al., 2023).
 However, none of these works consider model robustness under domain shift between training and testing data.
-Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains.
+Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022b) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains.
 2) The evaluations are restricted to the case when the number of clients is equal to the number of domains, which may be an unrealistic assumption (e.g., a hospital that has multiple imaging centers or a device that is used in multiple locations).
@@
 On the other hand, there are also benchmarks specifically designed for FL.
-For instance, the LEAF benchmark (Caldas et al., 2018) provides a standardized framework for evaluating FL algorithms.
-It includes several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario.
+For instance, the LEAF (Caldas et al., 2018) and FLamby (Terrail et al., 2022) benchmark provides a standardized framework for evaluating FL algorithms.
+They include several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario.
 Another benchmark for FL is PFLBench (Chen et al., 2022), which focuses on evaluating personalized FL methods.
@@
 Major contributions: We develop a benchmark methodology for evaluating Federated DG with various client heterogeneity contexts and diverse datasets, and we evaluate representative Federated DG approaches with this methodology.
-1) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 3).
-2) We propose a standardized definition of domain-based client heterogeneity that is unique to the FL context and interpolates between domain homogeneity and domain separation (see Fig.
-1) while controlling class imbalance.
-In particular, we develop a novel method to split any dataset with domain labels across any number of clients (see subsection 3.2).
-3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG.
+1) We develop a novel method to partition dataset across any number of clients that is able to control the heterogeneity among clients.
+(see Section 3).
+2) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 4).
+3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG on $7$ diverse datasets.
 Our results indicate that there still exist significant gaps and open research directions in Federated DG.
-4) We will release an extendable open-source library for evaluating Federated DG methods.
-Notation
+4) We release an extendable open-source library for evaluating Federated DG methods.
 […]
-2 Approaches to Federated Domain Generalization
-We first briefly review the DG problem, extend to the Federated DG problem, and explain domain-based client heterogeneity. Then, we discuss three categories of Federated DG approaches.
+2 Background: Federated Domain Generalization Methods
+In this section, we briefly introduce the problem backup and setup.
+Furthermore, evaluation on FL regimes often requires data partition methods to distribute training data to each client.
+Thus we also introduce the existing partition methods in this section.
 2.1 Problem Background and Setup
-Domain generalization (train-test heterogeneity)
-Unlike standard ML which assumes the train and test are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of the test domain distributions when only samples from the train domain distributions are given.
-Formally, given a set of train domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
+Domain generalization: Unlike standard ML which assumes the training and test data are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of test domain distributions using only samples from the training domain distributions.
+Formally, given a set of training domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
 $\min_{\theta}{\textstyle\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{d\in\mathcal{D}_{\text{test}}}}\mathcal{L}(\theta;p_{d})\quad\text{or}\quad\min_{\theta}\max_{d\in\mathcal{D}_{\text{test}}}\mathcal{L}(\theta;p_{d})\,,$ (1)
 where $\mathcal{L}(\theta;p_{d})=\mathbb{E}_{(x,y)\sim p_{d}}[\ell(x,y;\theta)]$ where $\ell$ is a per-sample loss function such as squared or cross-entropy loss.
-The key challenge in DG is that the train and test domain distributions are disjoint, i.e., $\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$, and thus, the method must be able to generalize beyond the train domain distributions to perform well on the test domain distributions.
-The naïve approach is to simply ignore the domains and perform empirical risk minimization (ERM) on all the training data—which is actually a challenging baseline to outperform in practice (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021).
-Federated DG
-Federated DG adds a layer of complexity because now the domain distribution samples are not centrally located.
-Instead, each FL client can update their local model based only on their local client distribution $p_{c}$ and then pass their local parameters $\theta_{c}$ to a central server, which will aggregate the local models and broadcast the model to all clients.
-The FL problem can be abstracted as follows:
-$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Optimize}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\quad\text{and}\quad\underbrace{\theta_{\text{global}}=\texttt{Aggregate}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
-where the client distributions may be homogeneous (i.e., $\forall(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Optimize minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$, and Aggregate aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
-Domain-based client heterogeneity
-While client heterogeneity (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client.
+Those two objectives are the ultimate goal for domain generalization.
+In the real-world setting, we never have access to $\mathcal{D}_{\text{test}}.$ There are works establishing objectives to approximate those two objectives.
+For example, Fish (Shi et al., 2021) is constructed using average training loss with a penalty, and IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2019) are inspired from the worst-case loss over a set of data distributions constructed from the training domains.
+There are mainly two kinds of DG test scenarios.
+The first scenario is the general DG where the test dataset form a new domain.
+Another scenario is called subpopulation shift where the test set is a subpopulation of the training distributions with the goal of better performance on the worst-case domain [^1] In Wilds (Koh et al., 2021), they treat subpopulation shift as another kind of train-test heterogeneity.
+DG is challenging where it breaks the i.i.d.
+assumptions in traditional machine learning.
+Federated DG: Federated DG represents an intuitive progression from centralized DG.
+Distributed learning is pivotal in numerous real-world applications, and its inherent architecture often results in multi-domain data.
+The quest for improved generalization is instinctive in this context.
+However, Federated DG introduces add two layers of complexity.
+First, the heterogeneous client distribution are harder to converge even without considering the train-test heterogeneity (Zhao et al., 2018; McMahan et al., 2017; Karimireddy et al., 2020; Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019).
+Second, privacy considerations and communication constraints typically prohibit the direct comparison of data across domains, thereby limiting the applicability of many conventional centralized DG methods.
+Formally, the FL problem can be abstracted as follows:
+$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Opt}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\;\text{and}\;\underbrace{\theta_{\text{global}}=\texttt{Agg}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
+where the client distributions may be homogeneous (i.e., $\forall\,(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists\,c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Opt minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$.
+The most common objective in Federated learning is empirical risk minimization (ERM), which minimizing the average loss over the given dataset.
+Agg aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
+Domain-based client heterogeneity: While previous client heterogeneity (i.e., $\exists\,c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client.
 At one extreme, FL with i.i.d.
 data would be equivalent to the mixture proportions being the same across all clients, i.e., $\forall c,c^{\prime},d,w_{c,d}=w_{c^{\prime},d}$, which we call the homogeneous setting.
-On the other extreme, if the number of clients and number of domains are equal ($D=C$), the domain weights could be disjoint between clients such that each client “owns” one or more domains, i.e., for any domain $d\in[D],$ $w_{c,d}>0\Leftrightarrow w_{c^{\prime},d}=0,\forall c^{\prime}\neq c$, which we call domain separation (in subsection 3.2, we extend the domain separation to the case $D\neq C$).
-Finally, we use $\lambda$ to denote an interpolation parameter between these two extremes: homogeneous case $\lambda=1$ and domain separation case $\lambda=0$.
-We give details of an explicit algorithm to use in practice for splitting datasets for evaluation in subsection 3.2 which covers two extremes and interpolation cases $\lambda\in(0,1)$.
+On the other extreme where the heterogeneity is maximum given the client number, we call complete heterogeneity (in Section 3.
 Table 1 summarizes the train-test heterogeneity in DG and the domain-based client heterogeneity from the FL context, where we focus on Federated DG.
-Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the train and test datasets (columns).
+Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the training and test datasets (columns).
 This paper focuses on the DG setting (right column).
@@
 Heterogeneous ($10<\lambda<$ 1) FL with heterogeneous clients Federated DG with heterogeneous clients
-Domain separation ($\lambda=0$) FL with domain separation Federated DG with domain separation
-Overview of Federated DG methods
-In this benchmark study, we explore three categories of Federated DG methods: DG methods originally designed for the centralized setting, FL methods specifically tailored to handle client heterogeneity, and methods specifically designed for Federated DG.
-To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories and compare to vanilla FedAvg (McMahan et al., 2017) with ERM loss, where any heterogeneity is simply ignored.
-These methods are selected based on their prominence and potential effectiveness for Federated DG.
-To ensure a diverse range of evaluation scenarios, we conduct experiments on various datasets and under various heterogeneity settings.
-Our results provide insights into their relative performance and suitability for different Federated DG contexts.
-This benchmark study aims to offer a broad overview of the current popular methods, enabling researchers and practitioners to make informed decisions when tackling Federated DG.
-2.2 Centralized DG methods adapted to the FL setting
-The first natural choice is directly migrating the DG methods from the centralized regime.
-To adapt those methods, we simply run the centralized DG method at each client locally with their own local dataset (see subsection 3.2 for how the local datasets are created), and then compute an average of model parameters at each communication round (see next paragraph).
-This approach is straightforward for the homogeneous ($\lambda=1$) and heterogeneous ($\lambda=0.1$) settings where each client has data from all training domains—albeit quite imbalanced for $\lambda=0.1$.
-This can be seen as biased updates at each client based on biased local data.
-In the domain separation case $\lambda=0$, (i.e., if all clients only have one primary domain, i.e., $\forall k,|P_{k}|=1$), this simple approach cannot be applied because centralized DG methods require data from at least two domains.
-In fact, these centralized DG methods degenerate to vanilla FedAvg if there is only one domain per client.
-Extending these methods to the case where all clients only have one domain without violating the FL constraints is an interesting direction for future work.
-A predominant and effective centralized DG approach is through representation learning, including domain-invariant representation learning by learning domain-invariant features via either kernel methods (Muandet et al., 2013; Ghifary et al., 2016) or invariant risk minimization (Arjovsky et al., 2019; Krueger et al., 2020), and domain adversarial neural networks Sun and Saenko (2016), and Li et al.
-(2018b) which explicitly tries to align the feature representation distribution.
-Besides invariant representation, there are other methods the general learning strategy to promote the generalization ability gradient operation Shi et al.
-(2021); Rame et al.
-(2022), which tries to learn generalized representations by directly operating on gradients.
-Other approaches include distributionally robust optimization Sagawa et al.
-(2019), which learns the worst-case distribution scenario of training domains; and meta-learning Finn et al.
-(2017); Li et al.
-(2018a), which is based on the learning-to-learn mechanism to learn general knowledge by constructing meta-learning tasks to simulate domain shift.
-We selected IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), MMD (Li et al., 2018b), Mixup (Zhang et al., 2017), DeepCoral (Sun and Saenko, 2016), and GroupDRO (Sagawa et al., 2019) from this category for their representative mechanisms.
-2.3 FL methods tackling client heterogeneity
-Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task).
-The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al.
+Complete Heterogeneity ($\lambda=0$) FL with complete heterogeneity Federated DG with complete heterogeneity
+2.2 Current Data partition methods
+In the FL context, the method for partitioning training data is crucial.
+As of now, there are three primary techniques for dataset partitioning.
+These methods are all fall in short in some properties we need.
+See Section 3 for a comprehensive analysis.
+In this section, we primarily present these methods as a background.
+Shards partitioning (McMahan et al., 2017) is a deterministic partition method.
+The datasets are initially ordered according to their labels.
+Then the datasets are partitioned into $2\times C$ shards, and each client $c\in[C]$ is allocated $2$ shards.
+Dirichlet Partitioning (Yurochkin et al., 2019; Hsu et al., 2019) is a stochastic partition method where each client hold the same number of samples but the proportion among labels are stochastic following a probability vector sampled from the Dirichlet distribution $\text{Dir}(\alpha p)$, where $p$ represents the prior label distribution on the whole training dataset, and $\alpha$ control the heterogeneity level.
+As $\alpha\rightarrow 0$, each client predominantly holds samples of a single label.
+Conversely, as $\alpha\rightarrow\infty$, the distribution of labels for each client becomes identical.
+Semantic Partitioning (Yuan et al., 2022) tries to create client heterogeneity on the features of data samples.
+Firstly, the data is processed through a pretrained model.
+The outputs from this model are then used to fit a Gaussian Mixture Model that features $K$ clusters for each class $Y$.
+These $KY$ clusters are then merged iteratively using an optimal bipartite for random label pairs.
+3 Heterogeneous Partitioning Method
+In this section, to thoroughly assess current Federated DG methods, we formulate an optimization problem that an ideal partition method should aim to solve.
+We then demonstrate that previous partition methods (subsection 2.2) are not feasible solutions.
+Subsequently, we introduce a new partition method, termed Heterogeneous Partitioning, which is an approximate optimal solution.
+Generally, our partition method effectively handles partitioning $D$ types of integer-numbered objects into $C$ groups.
+It’s broadly applicable, suitable for domain adaptation, ensuring fairness in Federated Learning (FL), and managing non-iid FL regimes.
+Table 2: Comparison of different partition methods.
+$C_{1}$ $C_{2}$
+Shards No Yes
+Dirichlet Yes No
+Semantic No Yes
+Ours Yes Yes
+Assume we have a set of all samples $\mathcal{S}=\{(x_{i},d_{i})\}_{i=1}^{n}$ and let $\mathcal{S}_{c}$ denote the set of samples assigned to the $c$-th client.
+For each client $c\in[C],$ define its domains as the set of domains from which $c$ has at least one sample., i.e., $\mathcal{D}_{c}=\{d\in[D]:\exists(x_{i},d_{i})\in\mathcal{S}_{c},d_{i}=d\}$.
+Denote $\mathcal{P}$ as a partition method.
+For Federated DG, it is common that clients have datasets collected from disjoint domains Koh et al.
 (2021).
-Several subsequent works have analyzed client heterogeneity by assuming bounded gradients (Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019) or bounded gradient dissimilarity (Li et al., 2020), and additionally assuming bounded Hessian dissimilarity (Karimireddy et al., 2020; Khaled et al., 2020; Liang et al., 2019).
-From this category, we selected FedProx (Li et al., 2020), which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model.
-Scaffold (Karimireddy et al., 2020) utilizes variance reduction to account for client heterogeneity.
-2.4 FL methods designed for Federated DG
-Limited research has focused explicitly on solving the Federated DG by design.
-FedDG Liu et al.
-(2021) introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol.
-Another approach, FedADG (Zhang et al., 2021), utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator.
-FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach.
-However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging.
-A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) (Tenison et al., 2022) aims to enhance generalization across clients and the global model.
-FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components.
-FedSR (Nguyen et al., 2022) proposes a simple algorithm that uses two locally-computable regularizers for domain generalization.
-Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
-3 Benchmark Methodology
-In this study, we aimed to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup.
-We evaluated a total of $13$ methods, encompassing three different types of approaches.
-3.1 Four evaluation dimensions
-(1) Domain-based client heterogeneity and (2) number of clients
-Previous studies on Federated DG have often focused on domain separation client heterogeneity where the number of clients equals the number of training domains, i.e., $C=D$.
-However, this excludes evaluation of the homogeneous and partially heterogeneous settings and restricts the number of clients to the number of training domains $D$.
-In particular, many pseudo-realistic datasets, such as those in DomainBed (Gulrajani and Lopez-Paz, 2020), which consist training domains $D\leq 5$, which limits the evaluation of methods under scenarios with a large number of clients.
-Conversely, when using realistic domain datasets that have many domains, such as those in WILDS (Koh et al., 2021), most current methods perform poorly under this extremely challenging setting.
-By introducing the domain split method discussed in subsection 3.2, we can explore various levels of client heterogeneity and relax the assumption that $C=D$ so that we can leverage both pseudo-realistic and realistic datasets and evaluate methods at an appropriate difficulty level.
-(3) Dataset difficulty and (4) dataset type
-While most Federated DG work focuses on standard image-based datasets, we evaluate methods across a broad range of dataset difficulty (ranging from easy pseudo-realistic datasets to very challenging realistic datasets) and two types of datasets ($3$ image datasets and $2$ text datasets).
-This ensures that we can fully understand the performance of each method across a wide range of scenarios.
-Further details can be found in Section subsection 3.3.
-3.2 Domain-based client heterogeneity by splitting DG datasets
-For evaluation dimensions (1) and (2) above, we need to control over the amount of client heterogeneity, ranging from homogeneous to domain separation, and we need to control over the number of clients, which may be smaller or larger than than the number of domains (i.e., $D\neq C$).
-We propose a way to split any DG dataset into any number of clients that allows explicit control of the amount of heterogeneity through the $\lambda$ hyperparameter which attempts to balance the number of samples per client.
-An illustration of our domain splitting procedure can be found in the following Fig.
+Therefore, it’s crucial that $\mathcal{P}$ can generate the strongest possible heterogeneity given the datasets and clients to reflect those common scenarios.
+This requirement can be encoded as either a single domain per client or no intersection of client domains, depending on the numbers of clients and domains (see constraint $C_{1}$).
+Furthermore, in distributed setting, to ensure a reasonable comparison with respect to domain generalization in the centralized setting, the total datasets received over the distributed system should match that of the centralized setting; and due to the privacy protocol, clients should not share the datasets.
+Thus, we need to preclude the partitions $\mathcal{P}$ where some data from $\mathcal{S}$ is not used by any client or clients either have duplicated data, which translate to constraint $C_{2}$.
+Thus, the feasible region of partition $\mathcal{P}$ is defined as $\mathcal{P}\in C_{1}\cap C_{2}.$
+Complete Heterogeneity (constraint $C_{1}$): $\mathcal{P}$ can generate the possibly strongest domain heterogeneity given the datasets and clients, i.e., $\mathcal{P}\in C_{1}$ where $C_{1}$ is defined as
+$C_{1}\triangleq\left\{\mathcal{P}\,\bigg{|}\,\text{ if $C>D,$ then $|\mathcal{D}_{c}|=1,\forall c\,;$ if $C\leq D,$ then $|\mathcal{D}_{c}\cap\mathcal{D}_{c^{\prime}}|=0,\forall c\neq c^{\prime}.$}\right\}$ (2)
+True Partition (constraint $C_{2}$): $\mathcal{P}$ must produce true set partitions $\{\mathcal{S}_{c}\},c\in[C]$.
+A true set partition is a non-empty, exhaustive, and pairwise disjoint family of sets, i.e., $\mathcal{P}\in C_{2}$ where $C_{2}$ is defined as
+$C_{2}\triangleq\left\{\mathcal{P}\,\bigg{|}\,\text{ $\mathcal{S}_{c}\neq\emptyset,\forall c$, $\bigcup_{c=1}^{C}\mathcal{S}_{c}=\mathcal{S}$, and $\mathcal{S}_{c}\cap\mathcal{S}_{c^{\prime}}=\emptyset,\forall c\neq c^{\prime}.$}\right\}.$ (3)
+[…]
+$\widehat{\mathcal{P}}\in\operatorname*{arg\,min}\limits_{\mathcal{P}\in C_{1}\cap C_{2}}\frac{1}{C-1}\sum_{c=1}^{C}\|n_{c}-\bar{n}\|_{2}^{2},$ (4)
+where $n_{c}=|\mathcal{S}_{c}|,$ $\bar{n}=\frac{1}{C}\sum_{c}n_{c}$.
+Notice that in many scenarios, it is not always able to achieve 00 loss, because $n_{c}\equiv\bar{n}$ may not be feasible for constraints $1$ and $2$ For example, consider the case of complete heterogeneity with $C=D$ but imbalance of domains, it is impossible to achieve balance $n_{c}\equiv\bar{n}$ and complete hetereogeneity $|\mathcal{D}_{c}|=1.$ Most previous partitions forces $n_{c}\equiv\bar{n}$ thus violating the constraints.
+In particular, Dirichlet Partitioning violates $C_{2}.$ Shard and semantic violate $C_{1}.$ We compare different partition methods in Table 2.
+See subsection A.2 for a detailed discussion on them.
+We construct Heterogeneous Partitioning $\{\mathcal{P}_{\lambda}\},\lambda\in[0,1],$ to solve Eqn.
+4.
+It is parameterized by $\lambda,$ where it can generate Complete Hetereogeneity ($\lambda=0$) and always satisfy True Partition (for all $\lambda\in[0,1]$) by construction; further it aims to achieve the best possible balance given these constraints.
+Heterogeneous Partitioning allows samples allocation from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via parameter $\lambda$, ranging from homogeneity to complete heterogeneity.
+In addition, when $\lambda=0,$ Heterogeneous Partitioning is optimal for Eqn.
+4 when there’s more clients than domains $C>D,$ Further, when $C\leq D,$ Eqn.
+4 is NP-hard problem proposed by Graham (1969) and Heterogeneous Partitioning is a linear time greedy approximation.
+The proof is deferred to subsection A.1.
+Heterogeneous Partitioning contains the following two steps, see Algorithm 1 for pseudo code.
+Step 1: Constructing complete heterogeneity.
+We explicitly construct $\mathcal{P}_{0}$ to satisfy $C_{1}$ while greedily trying to balance the variance.
+The key idea to create the complete heterogeneity is to partition one domain among the fewest possible clients.
+We adopt the following approaches depending on the size of the network $C$ and total domains $D$.
+Case I: If $C\leq D$, the domains are sorted in a descending order according to number of sample size in each domain, and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples, that is
+$c^{*}\triangleq\operatorname*{arg\,min}\limits_{c\in[C]}\sum_{d^{\prime}\in\mathcal{D}_{c}}n_{d^{\prime}},$ (5)
+where $n_{d}$ denotes the total sample size of domain $d.$ In this case, $\mathcal{P}_{0}$ ensures that no client shares domains with the others, i.e., $|D_{c}\cap D_{c}^{\prime}|=0,$ but attempts to balance the number of training samples between clients.
+Case II: If $C>D$, we first assign all the domains one by one to the first $D$ clients; then, starting from client $c=D+1$, we divide the currently largest average domain $d^{*}$ between the new client and the original clients associated with it.
+That is, for all $c\in\{D+1,D+2,\dots,C\},$ we assign $d^{*}$ to $c,$ and reassign the same sample sizes for those $c^{\prime}\in\{1,\dots,c\}$ which share the domain $d^{*}$ as
+$d^{*}\in\operatorname*{arg\,max}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]},\quad\frac{n_{d^{*}}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d^{*}\in\mathcal{D}_{c^{\prime}}]},$ (6)
+where rounding to integers is carefully handled when not perfectly divisible.
+Notice that in this case, some clients may share one domain, but no client holds two domains simultaneously, that is $|D_{c}|=1,$ for $c\in[C].$ In this case, $\mathcal{P}_{0}$ attempts to balance the number of samples across clients as much as possible while partitioning one domain among the fewest possible clients.
+Step 2: Computing domain sample counts for each domain with interpolation.
+We define $\mathcal{P}_{\lambda}$ through the sample counts for each client $c\in[C]$ per domain $d\in[D]$, denoted as $n_{d,c}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$
+$\mathcal{P}_{\lambda}:\quad n_{d,c}(\lambda)=\underbrace{\lambda\frac{n_{d}}{C}}_{\text{homogeneous}}+\underbrace{(1-\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}n_{d}}_{\text{complete heterogeneous}}\,,$ (7)
+where $\mathcal{D}_{c},c\in[C]$ are constructed by $\mathcal{P}_{0}$ Step $1$ This is simply a convex combination between homogeneous clients ($\lambda=1$) and extreme heterogeneous ($\lambda=0$).
+A smaller value of $\lambda$ corresponds to a more heterogeneous case.
+Given the number of samples per client $i\in[C]$ per domain $d\in[D]$, we simply sample $n_{d,c}$ without replacement from the corresponding domain datasets and build up the client datasets.
+Step $2$ ensures our partition $\mathcal{P}_{\lambda}$ satisfy the constraint $C_{2},$ because all samples are selected (as proven by $\sum_{c}n_{d,c}=n_{d})$ and there is no overlap (sample without replacement), and no client has zero samples given $|D_{c}|>0$ in the $\lambda=0,1$ cases.
+4 Benchmark Methodology and Evaluations
+In this section, we aim to conduct a comprehensive evaluation of the Federated DG task by considering four distinct dimensions of the problem setup equipped by Heterogeneous Partitioning.
+The four dimensions are 1) dataset type; 2) data difficulty; 3) domain-based client heterogeneity; and 4) the number of clients.
+This section is organized as follows: we introduce our benchmark datasets subsection 4.1 covering the first two dimensions, next we introduce $14$ methods we included in our evaluation from three different approaches subsection 4.2, then we introduce our benchmark setting and the evaluation results on all selected methods in subsection 4.3 brought the domain-based client heterogeneity and number of clients into consideration.
+4.1 Dataset Type and Dataset Difficulty Metrics
+While most Federated DG work focuses on standard image-based datasets, we evaluate methods over a diverse selection of datasets.
+These datasets encompass multi-domain datasets from simpler, pseudo-realistic ones to the considerably more challenging realistic ones.
+Specifically, our study includes five image datasets and two text datasets.
+Additionally, within these $7$ datasets, we include one subpopulation shift within image datasets (CelebA) and another within text datasets (Civilcomments).
+Furthermore, our dataset selections span a diverse range of subjects including general objects, wild camera traps, medical images, human faces, online comments, and programming codes.
+We also introduce dataset difficulty metrics to measure the empirical challenges of each dataset in the Federated DG task.
+Dataset Difficulty Metrics.
+To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities.
+We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context using the baseline objective ERM.
+For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance without and with samples from the test domain (i.e., the later is able to “cheat” by seeing part of test domain samples during training).
+For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples.
+These dataset difficulty metrics can be formalized as follows, for all $\lambda\in[0,1]$
+$R_{\text{DG}}\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})},R_{\text{FL}}(\lambda)\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
+where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the training domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ partition respectively of $\mathcal{S}_{\text{DG-test}}$.
+For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the training domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from the DG effect.
+We apply these metrics in our $7$ selected datasets and include the summary table in Table 6.
+Smaller $R_{\text{DG}}$ and $R_{\text{FL}}(\lambda)$ indicate a more challenging dataset.
+For instance, FEMNIST has $R_{\text{DG}}=1$ indicates the lack of domain heterogeneity.
+To contrast, IwildCam and CivilComments have small $R_{\text{FL}}$ showing the challenges from the FL side.
+4.2 Benchmark Methods
+In this benchmark study, we explore three categories of Federated DG methods.
+The first category is centralized DG methods adapted into FL regimes.
+The second category is FL methods tackling client heterogeneity and the third category is Federated DG methods.
+Please see Appendix B for a detailed discussion on current available methods to solve Federated DG.
+To provide a comprehensive evaluation, we assess the performance of several representative methods from each of these categories.
+specifically, we choose IRM (Arjovsky et al., 2019), Fish (Shi et al., 2021), Mixup (Zhang et al., 2017), MMD (Gretton et al., 2006), Coral (Sun and Saenko, 2016), GroupDRO (Sagawa et al., 2019) and adapted them into FL regime by applying them in every local clients and keeping the aggregation identical to FedAvg, i.e., weighted averaging aggregation.
+All those methods collapse to FedAdam if locally there is only one domain available.
+We consider $4$ methods.
+FedDG (Liu et al., 2021), FedADG (Zhang et al., 2021), FedSR (Nguyen et al., 2022) and FedGMA (Tenison et al., 2022a), which are naturally designed to solve Federated DG.
+We also consider FedProx (Li et al., 2020) and Scaffold (Karimireddy et al., 2020) from the Federated methods tackling client heterogeneity approach.
+All those methods will be compared to two baselines.
+ERM objective with Adam optimizer (Kingma and Ba, 2014) and its FL counterpart FedAdam (Reddi et al., 2020), which is a variant of FedAvg (McMahan et al., 2017).
+4.3 Main Results
+In this section, we present the performance results of $14$ representative methods, derived from distinct research areas, on $7$ diverse datasets.
+For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison.
+Client number Equipped with our Heterogeneous Partitioning, we can explore various levels of client heterogeneity and relax the assumption that $C=D$ so that we can leverage both pseudo-realistic and real-world datasets and evaluate methods on larger scale of FL clients.
+In particular, we set the number of clients to $100$
+Validation domain In DG task, we cannot access the test domain data.
+However, we are particularly concerned about the model performance outside the training domains, thus we preserve a small portion of the domains we can access as held-out validation domains, and the held-out validation domains are used for model selection and early stopping.
+Please see Appendix D for more detail.
+After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3, Table 4 and Table 5.
+See subsection D.4 for detailed hyperparameters choices.
+We make the following remarks on the main results from Table 3, Table 4 and Table 5.
+Because FEMNIST has low DG difficulty, we defer the results on FEMNIST in the Appendix (Table 11).
+Remark 4.1.
+FedAvg with an ERM objective is a strong baseline, especially on the general DG datasets.
+We observe FedAvg-ERM outperforms other methods multiple times.
+It still servers a strong baseline that is challenging to beat across general DG datasets, similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021).
+We recommend always including FedAvg as a baseline in all future evaluations.
+Remark 4.2.
+Most centralized DG methods degrade in the FL setting.
+For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3.
+Further, degradation can be seen in PACS and CelebA when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
+Remark 4.3.
+FL methods tackling client heterogeneity help convergence.
+Notably, IWildCam and CivilComments datasets bring greater challenges in the model convergence as Table 6.
+In this scenario, FL methods tackling the client heterogeneity are able to better tackle this challenge while other methods all failed in this context, see Table 3 and Table 5.
+Given the fact that this is a common challenge in Federated DG, there’s a need for future research to simultaneously enhance the performance with both client heterogeneity and train-test heterogeneity.
+Remark 4.4.
+Federated DG methods still in need.
+Upon evaluating current Federated DG methods on larger network scales, they all fail to outperform ERM.
+Among the four methods we evaluating, FedDG performs the best and could achieve higher accuracy on PACS with $\lambda=0$.
+However, FedDG requires sharing all local datasets amplitude information to the aggregation server, thus brings privacy concerns.
+FedADG methods contains tens of hyperparameters, and due to the nature of GAN (Goodfellow et al., 2020), it is challenging to optimize.
+FedSR and FedGMA also did not outperform ERM objective.
+It shows a clear demand for further improvement.
+Remark 4.5.
+Addressing the subpopulation shift could be an initial step.
+In the centralized setting, all evaluated methods surpassed the performance of ERM on two subpopulation shift datasets: CelebA and CivilComments.
+However, in a Federated context, only FedDG managed to outpace ERM on CelebA, and yet it couldn’t match the performance of centralized approaches.
+Further exploration in federated DG is in demand to bridge this gap.
+Remark 4.6.
+The performance of real-world data significantly degrades as $\lambda$ decreases.
+This can be seen from IWildCam and Py150 dataset at Table 3 and Table 5.
+While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG.
+We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
+Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
+“-” means the method is not applicable in that context.
+Bold is for best and italics is for second best in each column.
+We report the standard deviation among 3 runs.
+Please see the Table 8 in the appendix for higher precision report.
+PACS ($D=2$) IWildCam ($D=243$)
+$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
+(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
+B
+FedAvg-ERM $\bf 0.94\pm 0.01$ $\bf 0.95\pm 0.02$ $\mathbf{0.96\pm 0.02}$ $\mathbf{0.95\pm 0.02}$ $\mathit{0.34\pm 0.00}$ $\mathbf{0.30\pm 0.00}$ $0.25\pm 0.00$ $\mathbf{0.20\pm 0.01}$
+DG Adapted IRM $0.92\pm 0.01$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.32\pm 0.00$ $0.20\pm 0.01$ $0.18\pm 0.01$
+Fish $\mathit{0.94\pm 0.02}$ $0.65\pm 0.13$ $0.35\pm 0.11$ $\mathbf{0.35\pm 0.00}$ $0.18\pm 0.00$ $0.15\pm 0.01$
+Mixup $0.92\pm 0.01$ $0.89\pm 0.05$ $0.83\pm 0.03$ $0.33\pm 0.01$ $0.09\pm 0.01$ $0.07\pm 0.01$
+MMD $0.93\pm 0.02$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.32\pm 0.00$ $0.19\pm 0.01$ $0.16\pm 0.01$
+DeepCoral $0.93\pm 0.01$ $0.92\pm 0.04$ $0.86\pm 0.04$ $0.33\pm 0.01$ $0.19\pm 0.01$ $0.16\pm 0.01$
+GroupDRO $0.93\pm 0.01$ $\mathit{0.94\pm 0.03}$ $0.95\pm 0.02$ $0.21\pm 0.00$ $0.13\pm 0.01$ $0.20\pm 0.01$
+FL FedProx - $0.90\pm 0.03$ $0.89\pm 0.03$ $0.90\pm 0.02$ $0.25\pm 0.00$ $0.20\pm 0.00$ $0.17\pm 0.00$
+Scaffold - $0.90\pm 0.03$ $0.89\pm 0.02$ $0.90\pm 0.02$ $\mathit{0.28\pm 0.00}$ $\mathbf{0.26\pm 0.00}$ $\mathit{0.18\pm 0.01}$
+AFL - $0.93\pm 0.03$ $0.92\pm 0.04$ $0.91\pm 0.04$ $0.26\pm 0.01$ $0.16\pm 0.01$ $0.03\pm 0.00$
+FDG FedDG - $0.93\pm 0.02$ $\mathit{0.95\pm 0.02}$ $\mathit{0.94\pm 0.02}$ $0.27\pm 0.00$ $0.24\pm 0.00$ $0.17\pm 0.00$
+FedADG - $0.94\pm 0.01$ $0.94\pm 0.00$ $0.94\pm 0.01$ $0.26\pm 0.01$ $\mathit{0.25\pm 0.01}$ $0.16\pm 0.00$
+FedSR - $0.64\pm 0.16$ $0.55\pm 0.09$ $0.54\pm 0.09$ $0.18\pm 0.01$ $0.14\pm 0.01$ $0.09\pm 0.00$
+FedGMA - $0.75\pm 0.05$ $0.73\pm 0.09$ $0.72\pm 0.01$ $0.22\pm 0.00$ $0.15\pm 0.00$ $0.09\pm 0.00$
+Table 4: Test accuracy on CelebA and Camelyon17 dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
+“-” means the method is not applicable in that context.
+Bold is for best and italics is for second best in each column.
+We report the standard deviation among 3 runs.
+Please see the Table 9 in the appendix for higher precision report.
+CelebA ($D=2$) Camelyon17 ($D=3$)
+$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
+(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
+B
+FedAvg-ERM $0.77\pm 0.04$ $0.61\pm 0.02$ $0.52\pm 0.04$ $0.45\pm 0.02$ $0.90\pm 0.01$ $\mathit{0.95\pm 0.01}$ $\mathit{0.95\pm 0.00}$ $\bf 0.95\pm 0.00$
+DG Adapted IRM $\mathbf{0.89\pm 0.01}$ $0.71\pm 0.05$ $0.74\pm 0.01$ $0.91\pm 0.06$ $0.95\pm 0.00$ $0.95\pm 0.00$
+Fish $\mathit{0.88\pm 0.01}$ $0.14\pm 0.16$ $0.07\pm 0.05$ $0.92\pm 0.01$ $0.88\pm 0.01$ $0.88\pm 0.01$
+Mixup $0.29\pm 0.498$ $0.13\pm 0.22$ $0.13\pm 0.22$ $\mathit{0.94\pm 0.01}$ $\mathbf{0.95\pm 0.00}$ $0.95\pm 0.01$
+MMD $0.84\pm 0.05$ $0.81\pm 0.03$ $0.77\pm 0.03$ $0.92\pm 0.03$ $0.95\pm 0.00$ $0.95\pm 0.01$
+DeepCoral $0.85\pm 0.04$ $\mathbf{0.81\pm 0.02}$ $\mathbf{0.78\pm 0.02}$ $\mathbf{0.94\pm 0.01}$ $0.95\pm 0.00$ $0.95\pm 0.00$
+GroupDRO $0.87\pm 0.03$ $\mathit{0.81\pm 0.03}$ $0.76\pm 0.06$ $0.92\pm 0.03$ $0.95\pm 0.01$ $\mathbf{0.95\pm 0.00}$
+FL FedProx - $0.13\pm 0.24$ $0.12\pm 0.25$ $00.00\pm 0$ 0.00 - $0.94\pm 0.00$ $0.94\pm 0.00$ $\mathit{0.94\pm 0.01}$
+Scaffold - $0.73\pm 0.01$ $\mathit{0.78\pm 0.02}$ $\bf 0.80\pm 0.02$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.93\pm 0.00$
+AFL - $0.81\pm 0.02$ $0.83\pm 0.01$ $0.83\pm 0.01$ $0.94\pm 0.01$ $0.95\pm 0.01$ $0.93\pm 0.01$
+FDG FedDG - $0.56\pm 0.18$ $0.53\pm 0.11$ $0.46\pm 0.20$ $0.86\pm 0.01$ $0.86\pm 0.00$ $0.87\pm 0.02$
+FedADG - $0.67\pm 0.00$ $0.67\pm 0.01$ $\mathit{0.60\pm 0.01}$ $0.94\pm 0.00$ $0.93\pm 0.00$ $0.93\pm 0.00$
+FedSR - $0.38\pm 0.02$ $0.36\pm 0.04$ $0.38\pm 0.02$ $0.93\pm 0.01$ $0.92\pm 0.01$ $0.93\pm 0.00$
+FedGMA - $0.62\pm 0.01$ $0.62\pm 0.02$ $0.49\pm 0.01$ $0.90\pm 0.01$ $0.85\pm 0.01$ $0.82\pm 0.03$
+Table 5: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
+“-” means the method is not applicable in that context.
+Bold is for best and italics is for second best in each column.
+FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
+We report the standard deviation among 3 runs.
+Please see the Table 10 in the appendix for higher precision report.
+CivilComments ($D=16$) Py150 ($D=5477$)
+$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
+(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
+B
+FedAvg-ERM $0.54\pm 0.00$ $0.36\pm 0.03$ $0.35\pm 0.02$ $0.33\pm 0.01$ $\mathbf{0.68}\pm 0.00$ $\mathbf{0.68}\pm 0.00$ $\mathit{0.65\pm 0.00}$ $\mathit{0.64}\pm 0.00$
+DG Adapted IRM $0.64\pm 0.00$ $\mathit{0.59\pm 0.02}$ $0.53\pm 0.04$ $\mathit{0.68}\pm 0.00$ $\mathit{0.66\pm 0.00}$ $\mathbf{0.65\pm 0.00}$ $0.63\pm 0.00$
+Fish $\mathbf{0.67}\pm 0.00$ $0.42\pm 0.16$ $0.34\pm 0.17$ $0.66\pm 0.00$ $0.65\pm 0.00$ $0.65\pm 0.00$ $\mathbf{0.64\pm 0.00}$
+MMD $\mathit{0.65}\pm 0.00$ $\mathbf{0.63\pm 0.01}$ $\mathbf{0.61\pm 0.01}$ $0.66\pm 0.00$ $0.63\pm 0.00$ $0.63\pm 0.00$ $0.61\pm 0.00$
+DeepCoral $0.59\pm 0.00$ $0.52\pm 0.06$ $0.46\pm 0.08$ $0.66\pm 0.00$ $0.65\pm 0.00$ $0.65\pm 0.00$ $0.64\pm 0.00$
+GroupDRO $0.64\pm 0.00$ $0.48\pm 0.01$ $\mathit{0.47\pm 0.00}$ $0.51\pm 0.00$ $0.59\pm 0.00$ $0.60\pm 0.00$ $0.61\pm 0.00$
+FL FedProx - $0.18\pm 0.03$ $0.17\pm 0.03$ $0.17\pm 0.04$ $0.64\pm 0.00$ $0.63\pm 0.00$ $0.61\pm 0.00$
+Scaffold - $0.39\pm 0.02$ $0.38\pm 0.02$ $\mathit{0.33\pm 0.01}$ $0.64\pm 0.00$ $0.64\pm 0.00$ $0.62\pm 0.00$
+AFL - $0.55\pm 0.02$ $0.47\pm 0.01$ $\mathbf{0.44\pm 0.02}$ $0.49\pm 0.00$ $0.49\pm 0.00$ $0.47\pm 0.00$
+FDG FedSR - $0.36\pm 0.00$ $0.34\pm 0.00$ $0.32\pm 0.00$ $0.53\pm 0.00$ $0.53\pm 0.00$ $0.45\pm 0.00$
+FedGMA - $0.21\pm 0.03$ $0.20\pm 0.02$ $0.20\pm 0.02$ $0.62\pm 0.00$ $0.61\pm 0.00$ $0.60\pm 0.00$
+Additional DG challenges from FL.
+For further understanding, we explore some additional questions on several smaller datasets because it is computationally feasible to train many different models on these datasets.
+Specifically, we explore how the number of clients, amount of communication (i.e., the number of server aggregations) in the federated setup, and client heterogeneity affects the performance of various methods.
+The figures and detailed analysis are provided in the Appendix but we highlight two remarks here.
+Remark 4.7.
+The number of clients $C$ strongly influences performance.
+Performance in DG plunges from $90\%$ to as low as $10\%$ as $C$ shifts from $1$ to $200$ as suggested in Fig.
+2.
+We explore this on $4$ representative methods on three datasets.
+We advocate for future Federated DG assessments to consider larger number of clients, like $50$ to $100$ rather than only considering small numbers of clients.
+This demonstrates pressing, unaddressed challenges in Federated DG when grappling with many clients.
+Notably, FedADG and FedSR degrade faster with increasing number of clients.
+Remark 4.8.
+The number of communications does not monotonically affect the DG performance.
+In the FL context, we note a unique implicit regularization phenomenon: optimal performance is achieved in relatively few communication rounds Fig.
+3.
+For instance, with PACS, the ideal communication round is just $10$ (while maintaining constant computation).
+Contrarily, for in-domain tasks, FL theory indicates that increased communication enhances performance, as shown by (Stich, 2018, Theorem 5.1).
+Investigating how DG accuracy relates to communication rounds, and potential implicit regularization through early stopping, offers a compelling direction for future research.
+5 Conclusion and Discussion
+We’ve established a benchmark methodology for evaluating DG tasks in the FL regime, assessing 14 prominent methods on $7$ versatile datasets.
+Our findings indicate that Federated DG remains an unresolved challenge, with detailed gaps outlined in appendix Table 12.
+Recommendations for future evaluations of Federated DG.
+1) Evaluation on the dataset where $R_{\text{DG}}<1$.
+If $R_{\text{DG}}\approx 1$, there will be no real train-test heterogeneity.
+2) The FL regime introduces two main challenges for domain generalization (DG) methods.
+Firstly, the complete heterogeneity scenario limits the exchange of information between domains.
+Secondly, when ($R_{\text{FL}}<1$), it poses convergence challenges for the methods.
+One way to produce lower $R_{\text{FL}}$ is to increase the number of clients.
+In future evaluations, it would be valuable to assess the capabilities and limitations of proposed methods in handling these challenges.
+3) Always include FedAvg as the baseline comparison.
+4) Evaluation on both Subpopulation shift datasets and general DG datasets since the methods might perform differently.
+Suggestions for future work in Federated DG.
+1) The federated domain generalization (DG) task remains unsolved, and further work is needed to address the challenges.
+2) While our primary focus remains on domain generalization (DG), it is important for future methods to also address the issue of improving convergence when the FL learning rate ($R_{\text{FL}}$) is small.
+This consideration holds particular significance in real-world applications where efficient convergence is crucial.
+We hope this work provide a better foundation for future work in Federated DG and spur research progress.
+Reproducibility Statement
+Code for reproduce the results is available at the following link:
+https://github.com/inouye-lab/FedDG_Benchmark.
+We include detailed documentation in using our code to reproduce the results throughout the paper.
+We also provide documentation in adding new algorithm’s DG evaluation in the FL context.
+The code contains all the experiments included in section 4.
+To help reproducibility, we have
   • 1.
+    Include the requirements.txt file generated by conda to help environment setup.
+  • 2.
+    The code contains a detailed Readme file to help code reading, experiment reproducing and future new method implementing.
+  • 3.
+    We include the config file identical to the paper including the seed to help reproducibility.
+  • 4.
+    We include the command needed to run the experiments.
+  • 5.
+    The model structure and tokenizers are available.
+  • 6.
+    We include the hyperparameter grid search configuration to reproduce Table 7.
+  • 7.
+    The dataset preprocessing scripts, transform functions are also included.
+We hope our reproducible and extendable codebase could help both new methods implementations and better evaluation in the Federated DG fields.
+Acknowledgement
+This work was supported by Army Research Lab under Contract No.
+W911NF-2020-221.
+R.B.
+and D.I.
+also acknowledge support from NSF (IIS-2212097) and ONR (N00014-23-C-1016).
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
+Appendix
+[…]
+Appendix A Data Partition in Federated DG
+A.1 Heterogeneous Partitioning Algorithm and its guarantee
+Denote $\mathcal{P}_{0}$ as the complete heterogeneous case corresponding to Heterogeneous Partitioning with $\lambda=0.$ We have the following guarantees on the optimality of $\mathcal{P}_{0}.$
+Proposition A.1.
+When $C\geq D,$ $\mathcal{P}_{0}$ is optimal for Eqn. 4. When $C<D,$ Eqn. 4 is NP hard, and $\mathcal{P}_{0}$ is a fast greed approximation.
+Proof.
+Case I.
+When $C\geq D,$ $\mathcal{P}_{0}$ generates $\{S_{c}\},c\in[C]$ such that for all clients which share a same domain, their sample size is the same, that is, for all $c,c\in C^{\prime}$ share some domain $d\in D,$ there holds
+$n_{c}=n_{c^{\prime}}.$ (8)
+Further, for sample variance, there holds
+$\frac{1}{C-1}\sum_{c=1}^{C}\|n_{c}-\bar{n}\|_{2}^{2}$
+$=\frac{1}{2C(C-1)}\sum\limits_{i=1}^{C}\sum\limits_{j=1}^{C}\|n_{i}-\bar{n}-n_{j}+\bar{n}\|_{2}^{2}$
+$=\frac{1}{2C(C-1)}\sum\limits_{i=1}^{C}\sum\limits_{j=1}^{C}\|n_{i}-n_{j}\|_{2}^{2}.$ (9)
+Therefore, Eqn. 4 boils down to
+$\widehat{\mathcal{P}}\in\operatorname*{arg\,min}\limits_{\mathcal{P}\in C_{1}\cap C_{2}}\frac{1}{2C(C-1)}\sum\limits_{i=1}^{C}\sum\limits_{j=1}^{C}\|n_{i}-n_{j}\|_{2}^{2}.$ (10)
+Chaining with Eqn. 8, we have
+$\frac{1}{2C(C-1)}\sum\limits_{i=1}^{C}\sum\limits_{j=1}^{C}\|n_{i}-n_{j}\|_{2}^{2}=\frac{1}{2C(C-1)}\sum\limits_{i=1}^{D}\sum\limits_{j=1}^{D}\|\tilde{n}_{i}-\tilde{n}_{j}\|_{2}^{2},$ (11)
+where $\tilde{n}_{i}$ is the same sample size of clients which share the domain $i.$ The equality comes from the second equation in Eqn.
+6, where we reassign the same sample sizes for those $c^{\prime}\in\{1,\dots,c\}$ which share the domain $d^{*}$ as
+$\frac{n_{d^{*}}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d^{*}\in\mathcal{D}_{c^{\prime}}]}.$ (12)
+In addition, notice that we always divide the currently largest average domain $d^{*}$ between the new client and the original clients associated with it. That is,
+$d^{*}\in\operatorname*{arg\,max}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}.$ (13)
+Therefore, the discrepancy between $\tilde{n}_{i}$ and $\tilde{n}_{j},$ $i,j\in[D]$ is minimized.
+It is straightforward to verify such partition $\mathcal{P}_{0}$ also satisfy the true partition (constraint).
+Thus, $\mathcal{P}_{0}$ is optimal for Eqn.
+4.
+Case II. When $C<D$, the problem corresponds to the classic NP-Hard problem known as Multiway Number Partitioning. We employ a well-known linear-time greedy algorithm to address this [Graham, 1969].
 Figure 1: Distinct color refers to distinct domain data, and $\lambda$ is the domain balancing parameter.
 (a): train-test domain heterogeneity.
-(b): Domain split when $C\leq D$ and domain split when $C>D.$ (c): domain split when domain number exceeds client number, i.e., $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and separation ($\lambda=0$).
-See Algorithm 1 for domain split when $C>D.$
-Algorithm 1 DomainSplit Algorithm
+(b): domain partitioning when $C\leq D$ and domain partitioning when $C>D.$ (c): domain partitioning illustration when $C\leq D;$ homogeneous ($\lambda=1$), heterogeneous ($\lambda=0.1$), and extreme heterogeneous ($\lambda=0$).
+We provide the pseudo code for Heterogeneous Partitioning in Alg.
+1 and an illustration in Fig.
+1.
+In Fig.
+1, (a) represents a multi-domain dataset with $4$ training domains.
+Each color represents a domain and the height of the bar represents the number of samples in this domain.
+(b) shows the complete heterogeneity with two scenarios $C<D$ and $C>D$.
+(c) is a detailed partition result with different interpolating with the $C<D$ case.
+It can be seen that in the complete heterogeneity case, the partition satisfies constraint $1,2,$ further, the sample size between clients are roughly even.
+For other cases, True Partition (Constraint) is satisfied and also the sample size between clients are roughly even.
+A.2 Other Partition Methods
+Algorithm 1 Domain Partition Algorithm
 Input Number of clients $C$, heterogeneity parameter $\lambda\in[0,1]$, and samples from each domain $(\mathcal{S}_{1},\cdots,\mathcal{S}_{D})$, where $n_{1},\dots,n_{D}$ denote the number of samples per domain and w.l.o.g.
@@
      for $c=D+1,\dots,C$ do
-        // Find on average largest domain to split
+        // Find on average largest domain to partition
         $d^{*}\in\operatorname*{arg\,max}\limits_{d\in[D]}\frac{n_{d}}{{\textstyle\sum_{c^{\prime}=1}^{C}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]}$
@@
   end if
-  // Split samples across the clients
+  // Partition samples across the clients
   $\forall c,\mathcal{S}_{c}\leftarrow\emptyset$
@@
 Output $(\mathcal{S}_{1},\mathcal{S}_{2},\cdots,\mathcal{S}_{C})$
-In Algorithm 1, we provide a concrete algorithm for splitting samples from multiple domains across an arbitrary number of clients $C$ while controlling the amount of domain-based client heterogeneity via $\lambda$, ranging from homogeneous clients to domain separation.
-Our algorithm has two main steps.
-In Step 1, we assign “primary” domain indices $\mathcal{D}_{c}\subseteq[D]$ to each client $c\in[C]$ depending on $C$ and $D$.
-If $C\leq D$, the domains are sorted in descending order according to number of sample size and are iteratively assigned to the client $c^{*}$ which currently has the smallest number of training samples (denoted by $\sum_{d^{\prime}\in\mathcal{D}_{c^{*}}}n_{d^{\prime}}$).
-In this case, the algorithm ensures that no client shares domains with the others but otherwise attempts to balance the total number of training samples between clients.
-If $C>D$, we first assign the domains one by one to the first $D$ clients.
-Then, starting from client $c=D+1$, we iteratively assign the current on average largest domain $d^{*}$ to $c,$ accounting for the fact that the domain may already shared by different clients—notationally, on average is represented through dividing by $\sum_{c^{\prime}}\mathbbm{1}[d\in\mathcal{D}_{c^{\prime}}]$.
-In this case, some clients may share one domain, but no client holds two domains simultaneously while otherwise attempting to balance the number of samples across clients as much as possible.
-In Step 2, we define the sample counts for each client and domain, denoted $n_{c,d}(\lambda)$ based on the balancing parameter $\lambda\in[0,1]:$ $n_{d,c}(\lambda)=\lambda\frac{n_{d}}{C}+(1\!\!-\!\!\lambda)\frac{\mathbbm{1}[d\in\mathcal{D}_{c}]}{\sum_{c^{\prime}=1}^{C}\mathbbm{1}[d\in\mathcal{D}_{c}]}n_{d}\,,$ where rounding to integers is carefully handled when not perfectly divisible and where $\mathbbm{1}[\cdot]$ is the indicator function.
-This is simply a convex combination between homogeneous clients ($\lambda=1$) and domain separation ($\lambda=0$).
-Given the number of samples per client per domain, we simply sample without replacement from the corresponding domain datasets and build up the client datasets.
-3.3 Dataset Selection and Dataset Difficulty Metrics
-To ensure a comprehensive evaluation of current methods across different difficulty levels, we have curated a range of datasets with varying complexities.
-We define two dataset metrics to measure the dataset difficulty with respect to the DG task and with respect to the FL context.
-For DG difficulty, we compute $R_{\text{DG}},$ the ratio of the ERM performance with and without samples from the test domain (i.e., the former is able to “cheat” by seeing part of test domain samples during training).
-For FL difficulty, we attempt to isolate the FL effect by computing $R_{\text{FL}}(\lambda),$ the ratio of ERM-based FedAvg $\lambda$ client heterogeneity over centralized ERM on in-domain test samples.
-These dataset difficulty metrics can be formalized as follows:
-$R_{\text{DG}}$ $\triangleq\frac{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}^{\prime}_{\text{DG-test}})}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}}\cup{\mathcal{S}^{\prime\prime}_{\text{DG-test}}},\mathcal{S}^{\prime}_{\text{DG-test}})}$
-$R_{\text{FL}}(\lambda)$ $\triangleq\frac{\texttt{FedAvg\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}};\lambda)}{\texttt{ERM\_Perf}(\mathcal{S}_{\text{DG-train}},\mathcal{S}_{\text{IN-test}})},$
-where ERM_Perf is the performance of ERM using the first argument as training and the second for test, FedAvg_Perf is similar but with the client heterogeneity parameter $\lambda$, $\mathcal{S}_{\text{DG-train}}$ denotes samples from the train domains $\mathcal{D}_{\text{train}},$ $\mathcal{S}_{\text{DG-test}}$ denotes samples from the test domains $\mathcal{D}_{\text{test}},$ and $\mathcal{S}^{\prime}_{\text{DG-test}}$ and $\mathcal{S}^{\prime\prime}_{\text{DG-test}}$ are $20\%,80\%$ split respectively of $\mathcal{S}_{\text{DG-test}}$.
-For $R_{\text{FL}}(\lambda)$, we use $\mathcal{S}_{\text{IN-test}}$ (test samples from the train domains) instead of $\mathcal{S}_{\text{DG-test}}$ to isolate the FL effect from DG effect.
-We choose five datasets in our benchmark: FEMNIST from Caldas et al.
-(2018), PACS from Gulrajani and Lopez-Paz (2020), and IWildCam, CivilComments and Py150 from Koh et al.
-(2021).
-We summarize the statistics and difficulty metrics in Table 2 and provide the rationale of selecting these datasets in the Appendix.
-Table 2: Summary of selected datasets as well as their difficulty metric, where $n$ is the number of samples, $D$ is the number of domains, $C$ is the number of clients used.
-Dataset Statistics Difficulty Metric
-$n$ $D$ $C$ $R_{\text{DG}}$ $R_{\text{FL}}(\lambda)$
-$\lambda=1$ $\lambda=0.1$ $\lambda=0$
-FEMNIST $737036$ $2586$ $100$ $1.000$ $0.980$ $0.981$ $0.981$
-PACS $9991$ $2$ $100$ $0.960$ $1.000$ $1.000$ $1.000$
-IWildCam $203029$ $243$ $243$ $0.449$ $0.682$ $0.525$ $0.525$
-CivilComments $448000$ $16$ $100$ $0.984$ $0.618$ $0.533$ $0.532$
-Py150 $150000$ $8421$ $100$ $0.969$ $0.999$ $0.998$ $0.998$
-4 Benchmark Experimental Results
-In this section, we report the performance of $13$ representative methods from three lines of research on $5$ different datasets, where the FEMINIST results are provided in the appendix given the DG task simplicity (i.e., $R_{\text{DG}}\approx 1$).
-For each dataset, we fix the total computation and communication rounds for different methods for a fair comparison.
-We then select the model based on the held-out-domain validation set.
-After training, we choose the model according to the early-stopping at the communication round which achieves the best held-out-domain performance, and finally we evaluate the performance on the test-domain in Table 3 and Table 4.
-See Appendix for detailed hyperparameters choices.
-We also include the results according to in-domain validation early-stopping in the Appendix but the trends are similar.
-We make the following remarks on the main results from Table 3 and Table 4.
-Remark 4.1.
-FedAvg with an ERM objective is a strong baseline.
-Simple FedAvg with an ERM objective is a strong baseline that is challenging to beat across datasets (except for CivilComments), similar to the centralized case stated in DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021).
-We recommend always including FedAvg as a baseline in all future evaluations.
-Table 3: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
-“-” means the method is not applicable in that context.
-Bold is for best and italics is for second best in each column.
-PACS ($D=2$) IWildCam ($D=243$)
-$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
-(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
-B
-FedAvg-ERM $0.8379$ $0.8766$ $0.9144$ $\mathit{0.9377}$ $0.2727$ $0.1762$ $0.1241$ $0.0707$
-DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
-Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
-Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
-MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $\mathit{0.3134}$ $0.0718$ $0.0924$
-DeepCoral $\bf 0.9216$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
-GroupDRO $0.9060$ $\bf 0.9395$ $\bf 0.9437$ $0.2531$ $0.0530$ $0.0908$
-FL FedProx - $0.8826$ $0.8407$ $0.8892$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
-Scaffold - $0.8215$ $0.8796$ $0.9239$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
-FDG FedDG - $\mathit{0.9234}$ $\mathit{0.9275}$ $\mathbf{0.9521}$ $0.1445$ $0.1191$ $0.0503$
-FedADG - $0.1892$ $0.0592$ $0.0598$ $0.0049$ $0.0049$ $0.0049$
-FedSR - $0.1246$ $0.1263$ $0.1257$ $0.0056$ $0.0056$ $0.0056$
-FedGMA - $0.8820$ $0.8467$ $0.8446$ $0.0106$ $0.0957$ $0.0106$
-Table 4: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
-“-” means the method is not applicable in that context.
-Bold is for best and italics is for second best in each column.
-FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
-CivilComments ($D=16$) Py150 ($D=5477$)
-$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
-(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
-B
-FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $\mathit{0.6405}$
-DG Adapted IRM $0.6408$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
-Fish $\mathbf{0.6713}$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
-MMD $\mathit{0.6520}$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $0.6565$
-DeepCoral $0.5847$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
-GroupDRO $0.6383$ $0.6028$ $0.4954$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
-FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
-Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
-FDG FedSR - 0.3603 0.3389 0.3188 - 0.0915 0.0860 0.0587
-FedGMA - 0.3341 0.3157 0.3195 - 0.6628 0.3737 0.3576
-Remark 4.2.
-Most centralized DG methods degrade in the FL setting.
-For image datasets, the DG methods adapted to the FL setting (except for GroupDRO) show significant degradation in performance compared to their centralized counterparts as can be seen when comparing the $C=1$ column to the $C>1$ columns in Table 3.
-Further, degradation can be seen in PACS when moving from the homogeneous client setting ($\lambda=1$) to the heterogeneous client setting ($\lambda=0.1$).
-Remark 4.3.
-FL methods tackling client heterogeneity perform surprisingly well on the Federated DG task.
-FedProx and Scaffold, which were designed for client heterogeneity but not specifically for the DG task, perform quite well in the Federated DG setting and even perform the best for IWildCam and Py150.
-Thus, we suggest including these methods in future Federated DG evaluations and suggest that future works could focus on combining the strengths of these methods and DG methods.
-Remark 4.4.
-The performance of real-world data significantly degrades as $\lambda$ decreases.
-This can be seen from IWildCam and Py150.
-While it is challenging and expensive to run models for IWildCam and Py150, they show the largest differences between methods and demonstrates the real-world challenge of Federated DG.
-We suggest including IWildCam and Py150 in most future DG evaluations given their unique nature across datasets.
-Remark 4.5.
-A diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG.
-This conclusion is inspired from the following two observations.
-1) When comparing FL methods, we notice opposite trends between PACS and all other datasets (IWildCam, CivilComments, Py150) when increasing client heterogeneity based on $\lambda$.
-For PACS, the FL methods (except for FedADG and FedSR) surprisingly seem to improve when increasing client heterogeneity (i.e., $\lambda=0$).
-However, for other datasets, the expected trend exists such that increasing client heterogeneity produces worse performance.
-From this, we recommend that using PACS alone might be a misleading or at least incomplete evaluation of Federated DG methods.
-2) Centralized DG methods adapted to FL perform the best on CivilComments.
-This unexpected result suggests that DG methods in the centralized regime may be able to better accommodate subpopulation shift, which is the special kind of shift that is exhibited in CivilComments.
-Both of these observations emphasize that a diversity of evaluation datasets is important for holistically evaluating new methods for Federated DG tasks.
-Additional DG challenges from FL.
-For further understanding, we explore some additional questions on the PACS dataset because it is the most common and computationally feasible to train many different models.
-Specifically, we explore how the number of clients, amount of communication (i.e., the number of server aggregations) in the federated setup, and client heterogeneity affects the performance of various methods.
-The figures and detailed analysis are provided in the Appendix but we highlight two remarks here.
-Remark 4.6.
-The number of clients $C$ strongly affects on the overall performance.
-The DG performance drops from $90\%$ to $50\%$ or even 10% when varying $C$ from $1$ to $200$.
-We strongly recommend future Federated DG evaluations to consider larger number of clients such as $50$ to $100$ rather than only a very small number of clients.
-It indicates that there exist significant unresolved challenges in the domain of federated domain generalization when dealing with a large number of clients, which have not been adequately explored.
-In particular, FedADG and FedSR seem to be sensitive to the number of clients because they perform poorly when $C\geq 10$, while in the original papers they were only evaluated on a few clients $C=3$ (and we reproduce the good results with such $C$ in the Appendix).
-Remark 4.7.
-The number of communications does not monotonically affect DG performance.
-We observe an interesting implicit regularization phenomena in the FL context: different methods achieve their best performance when the communication rounds is relatively low.
-In particular, for PACS, this communication round is $10$ (while fixing the total amount of computations).
-This is surprising as in contrast for in-domain task, FL suggests more communications leads to better performance (Stich, 2018, Theorem 5.1).
-Further theoretical investigation of the dependence of DG accuracy on the communications rounds, and its possible relation to implicit regularization via early stopping is an interesting area for future work.
-5 Conclusion and Discussion
-We first build a systematic benchmark methodology for evaluating Federated DG that includes novel methodologies for splitting the data across clients and evaluation of dataset difficulty.
-We then evaluate $13$ representative methods from three relevant lines of research.
-Our evaluation shows that Federated DG is still unsolved and significant gaps remain between centralized DG performance and Federated DG performance.
-Therefore, Federated DG is ripe for additional research.
-Based on our evaluation and observation, here are some recommendations and suggestions for future work in Federated DG.
-Recommendations for future evaluations of Federated DG.
+Shards.
+By adjusting the constant from $2$ to other value, it’s possible to regulate the heterogeneity to a certain extent.
+However, if the dataset contains imbalance number of samples per domain, shards method could not achieve the complete heterogeneity.
+For instance, if a dataset contains $2$ domains with $20,30$ data samples respectively.
+Shard partitioning will not be able to assign each client with data from only one domain.
+Dirichlet Partitioning.
+It draws $q_{c}\sim\text{Dir}(\alpha p),$ where $p$ characterizes a prior domain distribution over $D$ domains for each client $c\in[C].$ $\alpha>0$ is a concentration parameter controlling the heterogeneity among clients.
+As $\alpha\rightarrow 0$, each client predominantly holds samples of a single label.
+Conversely, as $\alpha\rightarrow\infty$, the distribution of labels for each client becomes identical.
+This method is not a desirable partitioning in the following sense.
+1) If sampling with replacement, then the datasets across clients may not be disjoint, and some data samples may not be allocated to any client, therefore breaking properties $3$ and $4$ 2) If sampling without replacement, the method is inapplicable when the available data from one domain doesn’t meet the collective demand for this domain’s data from all clients, breaking property $2$
+Semantic Partitioning.
+This method aims to partition dataset based on the feature similarity without domain labels.
+It first processes the data using a powerful pretrained model.
+The outputs before the fully connected layer are later used as features.
+Then, thses features are fit a Gaussian Mixture Model to form $K$ for each clusters for each class $Y$.
+These $KY$ clusters are then merged iteratively using an optimal bipartite for random label pairs across different class.
+This partition method is more heuristic and highly rely on the quality of the pretrained model.
+Furthermore, since there is no hard code label, and the partition is based on KL divergence, this method could not give a complete heterogeneity which violate property $1$.
+Despite that this method could not guarantee property $1$, it is a good partition method to create “domain"-heterogeneity without having domain labels.
+Example 1.
+We provide a toy example of partition a dataset containing $5$ training domains with $10,20,30,40,100$ data samples respectively.
+We illustrate how shards and Dirichlet methods work and highlight that disjointedness flexibility and controllability are violated when our goal is to partition the data into $2$ clients.
+Shards partitioning might allocation $[10,20,20,0,50]$ samples to the first client, and $[0,0,10,40,50]$ samples to the second client.
+This lead to the share of domain $3$ and $5$ between the two clients showing the violation of Complete Heterogeneity (Constraint).
+The Dirichlet partitioning will sample a distribution from a Dirichlet distribution with $\alpha\approx 0$.
+However, no matter what distribution are sampled, at least $3$ domains will be neglected, leading to the violation of True Partition (Constraint).
+In contrast, our Heterogeneous Partitioning will allocate $[10,20,30,40,0]$ to the first client and $[0,0,0,0,100]$ to the second client, effectively maintaining Complete Heterogeneity (Constraint), True Partition (Constraint), in addition, the sample size difference is 00 in this case.
+Appendix B Current Methods in solving DG
+Centralized DG methods
+Most work solving in DG lies in the centralize regime.
+A predominant and effective centralized DG approach is through representation learning.
+Arjovsky et al.
+[2019] tries to learn domain-invariant feature by aligning the conditional distribution of $p(Y|X)$ among different domains.
+Sun and Saenko [2016], and Li et al.
+[2018] tries to explicitly align the first order and second order momentum in the feature space.
+There are also methods trying to promote the out-of-distribution generalization by posting constraint on the gradient information among different domains, where Shi et al.
+[2021] tries to align the gradient among different domains, and Rame et al.
+[2022] enforces domain invariance in the space of the gradients of the loss.
+Other approaches include distributionally robust optimization [Sagawa et al., 2019], where this method learns the worst-case distribution scenario of training domains.
+Furthermore, Zhang et al.
+[2017] is not specifically design for solving domain generalization, but found to be effective as well when we extrapolate the data from different domains in the training dataset.
+Xu et al.
+[2020] introduces domain-mixup which extrapolates examples and features across training domains.
+Federated DG methods
+Limited research has focused explicitly on solving the Federated DG by design.
+FedDG Liu et al.
+[2021] introduced a specific FL paradigm for medical image classification, which involves sharing the amplitude spectrum of images among local clients, violating the privacy protocol.
+Another approach, FedADG [Zhang et al., 2021], utilizes a generative adversarial network (GAN) framework in FL, where each client contains four models: a featurizer, a classifier, a generator, and a discriminator.
+FedADG first trains the featurizer and classifier using empirical loss and then trains the generator and discriminator using the GAN approach.
+However, this method requires training four models simultaneously and tuning numerous hyperparameters, making convergence challenging.
+A novel aggregation method called Federated Gradient Masking Averaging (FedGMA) [Tenison et al., 2022b] aims to enhance generalization across clients and the global model.
+FedGMA prioritizes gradient components aligned with the dominant direction across clients while assigning less importance to inconsistent components.
+FedSR [Nguyen et al., 2022] proposes a simple algorithm that uses two locally-computable regularizers for domain generalization.
+Given the limited literature on solving domain generalization (DG) in the federated learning (FL) setting, we selected all the aforementioned algorithms.
+Algorithm 2 Modified AFL
+Initialize $\theta$, $\beta$
+  Transmit $\theta$ to all clients $c\in[C].$
+  for $t=1,2,\dots,T$ do
+     // ——–Update $\theta$——–
+     Transmit $\beta$ to all clients $c.$
+     for $c=1,2,\dots,C$ do
+        Update $\theta_{c}$ using SGD with multiple iteration.
+        Submit $\theta_{c}$ to the server.
+     end for
+     $\theta\leftarrow\sum_{c}\frac{1}{n_{c}}\theta_{c}.$  // Weighted average.
+     // ——–Update $\beta$——–
+     Transmit $\theta$ to all clients $c$
+     for $c=1,2,\dots,C$ do
+        Calculate $\ell_{d,c}=\nabla_{\beta_{d}}\mathcal{L}_{c}$
+        Submit updated $\ell_{d,c}$.
+     end for
+     for $d=1,2,\dots,D$ do
+        $\beta_{d}\leftarrow\text{Proj}(\beta_{d}+\gamma_{\beta}\sum_{c}\ell_{d,c}).$
+     end for
+  end for
+Output $\theta$
+FL methods tackling client heterogeneity
+Another line of research in FL aims to guarantee convergence even under client heterogeneity, but these FL-based methods still assume the train and test datasets do not shift (i.e., they do not explicitly tackle train-test heterogeneity of the domain generalization task).
+We include this line of work because methods considering converging over a heterogeneous dataset might bring better DG ability implicitly.
+The empirical observation of the statistical challenge in federated learning when local data is non-IID was first made by Zhao et al.
+[2018].
+Several subsequent works have analyzed client heterogeneity by assuming bounded gradients [Yu et al., 2019, Basu et al., 2019, Wang et al., 2019, Li et al., 2019] or bounded gradient dissimilarity [Li et al., 2020], and additionally assuming bounded Hessian dissimilarity [Karimireddy et al., 2020, Khaled et al., 2020, Liang et al., 2019].
+From this category, we selected FedProx [Li et al., 2020], which addresses statistical heterogeneity by adding a proximal term to the local subproblem, constraining local updates to be closer to the initial global model.
+Scaffold [Karimireddy et al., 2020] utilizes variance reduction to account for client heterogeneity.
+FL methods tackling fairness
+Agnostic Federated Learning (AFL) [Mohri et al., 2019, Du et al., 2021] shares similarities with Domain Generalization in a Federated context.
+This is evident as both approaches address scenarios where the test distribution diverges from the training distribution.
+AFL constructs a framework that naturally yields a notion of fairness, where the centralized model is optimized for any target distribution formed by a mixture of the client distributions.
+Thus, AFL is a good method to evaluate especially when tackling subpopulation shift tasks.
+AFL introduces a minimax objective which is identical to the GroupDRO [Sagawa et al., 2019],
+$\min_{\theta}\max_{\beta\in\Delta_{D}}\mathcal{L}(\theta,\beta)=\sum_{d=1}^{D}\beta_{d}\ell_{d}(\theta),$ (14)
+where $\theta$ denotes the model parameter and $\beta$ is a weight vector taking values in the simplex $\Delta_{D}$ of dimension $D.$ AFL introduces an algorithm applying projected gradient ascent on $\beta$ and gradient descent on $\theta$.
+It is designed for $C=D,$ where they assumes each client as a domain.
+Further, it requires communication per iteration.
+We made the following two modifications to accommodate the general $C\neq D$ cases and expensive communication cost:
   1. 1.
-    Stronger Baselines - FedAvg should always be included because it is a strong baseline (Remark 4.1).
-FL methods designed for client heterogeneity (though not necessarily DG) should also be included given their strong performance (Remark 4.3).
+    To allow $C\neq D,$ we construct new objective as the following:
+    $\min_{\theta}\max_{\beta\in\Delta_{D}}\mathcal{L}(\theta,\beta)=\sum_{c=1}^{C}\mathcal{L}_{c}(\theta,\beta)=\sum_{c=1}^{C}\sum_{d=1}^{D}\beta_{d}\ell_{d,c}(\theta).$ (15)
   2. 2.
-    Realistic Datasets - Federated DG methods should be evaluated on more realistic datasets.
-In particular, FEMNIST and PACS behave quite differently with respect to client heterogeneity than the more realistic WILDS datasets (Remark 4.4).
-Thus, we recommend including both IWildCam and Py150 in future evaluations of Federated DG.
-CivilComments may be useful for evaluating realistic subpopulation shift (Remark 4.5).
-  3. 3.
-    Large Number of Heterogeneous Clients - Evaluations should include scenarios with a large number of clients with varying degrees of heterogeneity.
-A large number of clients poses unique challenges for most methods (Remark 4.6).
-Additionally, the heterogeneity of clients is more realistic and plays a significant role in evaluation (Remark 4.4).
-Suggestions for future work in Federated DG.
-  1. 1.
-    Handle Domain Separation Case - The domain separation scenario ($\lambda=0$) limits the exchange of information between domains, i.e., a client may only have data from a single domain.
-Centralized approaches cannot be adapted to this setting and current methods still struggle under this realistic client heterogeneity setting.
-  2. 2.
-    Increase Convergence Rate - In most benchmark experiments (except for PACS, which is an easy dataset), we have observed slow convergence, particularly on challenging real-world datasets such as IWildCam and Py150.
-Moreover, we have noticed even slower convergence when client heterogeneity is high (i.e., when $\lambda$ is small).
-In addition, we have found that federated methods designed to improve convergence actually performed quite well even though they were not designed for the DG task (Remark 4.3).
-Therefore, it is crucial for future methods to improve the convergence rate of Federated DG approaches.
-  3. 3.
-    Investigate the Effect of Communication Frequency - Because increased communication frequency may actually hurt DG performance (Remark 4.7), further investigation into the regularization effect of early stopping and infrequent communications in FedAvg may yield important insights.
-  4. 4.
-    Understand the Effect of the Number of Clients - Given that the performance of some methods degrade very quickly when increasing the number of clients beyond a few (Remark 4.6), the community could benefit from an improved theoretic and empirical understanding of how the number of clients affects Federated DG in terms of convergence, computational requirements, sample complexity, and DG performance.
-We hope this work provides a better foundation for future work in Federated DG and accelerates research progress.
-Acknowledgements
-This work was supported by Army Research Lab under Contract No.
-W911NF-2020-221.
-R.B.
-and D.I.
-also acknowledge support from NSF (IIS-2212097) and ONR (N00014-23-C-1016).
-Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor(s).
-Appendix
-[…]
-Reproducibility Statement
-Code for reproduce the results is available at the following link:
-https://github.com/inouye-lab/FedDG_Benchmark.
-We include detailed documentation in using our code to reproduce the results throughout the paper. We also provide documentation in adding new algorithm’s DG evaluation in the FL context.
-Appendix A Datasets and Difficulty Metric
-A.1 Dataset Introduction
-In this section, we introduce the datasets we used in our experiments, and the split method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
+    To reduce communication cost, we allow $\theta$ to be updated multiple iterations locally per communication.
+Further, we maintain the update of $\beta$ on the central server, as in AFL, given that it requires projection of the global updates onto the simplex.
+This projection is not equivalent to averaging the locally projected updates of $\beta.$
+Modified AFL avoids the frequent communication compared to AFL.
+Appendix C Datasets and Difficulty Metric
+C.1 Dataset Introduction
+In this section, we introduce the datasets we used in our experiments, and the partition method we used to build heterogeneous datasets in the training and testing phase as well as the heterogeneous local training datasets among clients in the FL.
 FEMNIST
@@
 It contains 243 training domains, 32 validation and 48 test domains.
-Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference Koh et al.
-[2021].
+Usually people cares about rare speicies, thus we utilize the macro F1 score as the evaluation metric instead of standard accuracy, as recommended in the original dataset’s reference [Koh et al., 2021].
 The $R_{\text{DG}}=0.449$ makes it a very challenging dataset for domain generalization.
+CelebA
+CelebA (Celebrity Attribute) [Liu et al., 2015] is a large-scale face attributes dataset.
+It’s one of the most popular datasets used in the facial recognition and facial attribute recognition domains.
+We use a variant of the original CelebA dataset from [Sagawa et al., 2019] using hair color as the classification target and gender as domain labels.
+This forms a subpopulation shift task which is a special kind of domain generalization.
+Camelyon17
+It is a medical dataset consisting of a set of whole-slide images from multiple institutions (as domain label) to detect breast cancer.
+It consist $3$ training domains, $1$ validation domain and $1$ test domain.
+We partition the data following wilds benchmark [Koh et al., 2021].
 CivilComments
 It is a real-world binary classification text-based dataset formed from the comments from different demographic groups of people, containing $8$ demographic group.
-The goal is to judge whether the text is malicious or not.
+The goal is to judge whether the text is malicious or not, which is also a subpopulation shift dataset.
 Py150
@@
 We evaluate models by the accuracy on the class and method tokens.
-A.2 Dataset Split Setup
-For each dataset, we first split the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset.
+Table 6: Summary of selected datasets as well as their difficulty metric, where $n$ is the number of samples, $D$ is the number of domains, $C$ is the number of clients used.
+Dataset Statistics Difficulty Metric
+$n$ $D$ $C$ $R_{\text{DG}}$ $R_{\text{FL}}(\lambda)$
+$\lambda=1$ $\lambda=0.1$ $\lambda=0$
+FEMNIST $737036$ $2586$ $100$ $1.000$ $0.980$ $0.981$ $0.981$
+PACS $9991$ $2$ $100$ $0.960$ $1.000$ $1.000$ $1.000$
+IWildCam $203029$ $243$ $100$ $0.449$ $0.869$ $0.714$ $0.571$
+CelebA $162770$ $4$ $100$ $0.661$ $0.760$ $0.805$ $0.797$
+Camelyon17 $410359$ $3$ $100$ $0.969$ $1.000$ $1.000$ $1.000$
+CivilComments $448000$ $16$ $100$ $0.984$ $0.618$ $0.533$ $0.532$
+Py150 $150000$ $8421$ $100$ $0.969$ $0.999$ $0.998$ $0.998$
+C.2 Dataset partition Setup
+For each dataset, we first partition the dataset into $5$ categories, namely training dataset, in-domain validation dataset, in-domain test dataset, held-out validation dataset and test domain dataset.
 For FEMNIST and PACS dataset, we use Cartoon and Sketch as training domain, Art-painting as held-out domain, Painting as test domain.
-For training domain, we split $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively.
-For IWildCam, CivilComments and Py150, we directly apply the Wilds official splits.
-Appendix B Benchmark Experimental Setting
-B.1 Model Structure
+For training domain, we partition $10\%,10\%$ of the total training domain datasets as in-domain validation dataset and in-domain test dataset respectively.
+For IWildCam, CivilComments and Py150, we directly apply the Wilds official partitions.
+Appendix D Benchmark Experimental Setting
+D.1 Model Structure
 In this benchmark, for image-based datasets, we use ResNet-50 He et al.
@@
 [2020] and CodeGPT Radford et al.
-[2019] respectively as recommended by Wilds.
-B.2 Model Selection
+[2019] respectively as recommended by Wilds [Koh et al., 2021].
+D.2 Model Selection
 We conduct held-out domain model selection with $4$ runs for each methods.
@@
 The results are reported based on the best run.
-B.3 Early Stopping
+D.3 Early Stopping
 We conduct early stopping using the held-out validation dataset in our evaluation.
 For each dataset and method, We first run certain communication rounds, then we select the model parameters which achieves the best performance on the validation set.
-We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix D.
-B.4 Hyperparameters
+We report the held-out validation dataset in the main paper, and we report the results using the in-domain validation set in Appendix F.
+D.4 Hyperparameters
 In this section, we present the hyperparameters selected for the evaluation.
-We opted to communicate per epoch for all experiments.
-For PACS, we run totally $50$ communication rounds.
-For IWildCam, we run $50$ communication rounds.
-For CivilComments, we run $10$ communication rounds.
-For Py150, we run $10$ communication rounds, and for FEMNIST, we run $40$ communication rounds.
-Please refer to the Table 5 to review other hyperparameters.
-Table 5: Hyperparameters used in the evaluation. c represents centralized learning regime, and f represents the federated learning regime
-PACS IWildCam CivilComments Py150 FEMNIST
-c f c f c f c f c f
-Fed Avg Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-IRM Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-Penalty weight $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $100$
-Penalty anneal iters $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$ $500$
-Fish Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-meta learning rate $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$ $0.5$
-Mixup Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-$\alpha$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$
-MMD Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-Penalty Weight $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
-Deep Coral Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-Penalty Weight $1$ $1$ $10$ $10$ $1$ $1$ $1$ $1$ $1$ $1$
-Group DRO Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-$\eta$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$
-Fed Prox Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-$\mu$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
-Scaf fold Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-FedDG Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning Rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
-FedADG Classifier optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGDs
-Classifier learning rate $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$ $0.0002$
-Gen optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
-Gen learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
-Disc optimizer SGD SGD SGD SGD SGD SGD SGD SGD SGD SGD
-Disc learning rate $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$ $0.0007$
-$\alpha$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$ $0.15$
-FedSR Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-$\ell_{2}$ regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.001$ $0.001$ $0.01$ $0.01$
-cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
-FedGMA Optimizer Adam Adam Adam Adam Adam Adam AdamW AdamW Adam Adam
-Learning rate $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$ $8\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$
-Server step size $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
-Mask threshold $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
-Appendix C Additional FL-specific challenges for domain generalization
-As mentioned in Section 4, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
-i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $K$ increases on PACS.
+We grid search 8 times of run starting with learning rate same as ERM and other hyperparameters from the methods’ original parameters.
+We than select the hyperparameter based on the best performance on the held-out domain validation set.
+Please refer to the Table 7 to review all hyperparameters.
+Table 7: Hyperparameters used in the evaluation.
+Dataset Name PACS CelebA Camelyon17 FEMNIST IWildCam CivilComments Py150
+Total Communication $80$ $20$ $20$ $40$ $50$ $10$ $10$
+ERM lr $3\times 10^{-5}$ $3\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-5}$
+IRM lr $3\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-6}$
+penalty weight $1000$ $1000$ $1000$ $100$ $100$ $10$ $1000$
+Fish lr $3\times 10^{-5}$ $1\times 10^{-3}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-6}$
+meta-lr $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ $1$
+Mixup lr $3\times 10^{-5}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $3\times 10^{-5}$
+$\alpha$ $0.2$ $0.2$ $0.1$ $0.2$ $0.2$
+MMD lr $3\times 10^{-5}$ $0.0001$ $1\times 10^{-3}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $1\times 10^{-5}$
+penalty weight $1$ $1$ $0.1$ $1$ $1$ $10$ $1$
+Coral lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-6}$ $5\times 10^{-5}$
+penalty weight $1$ $1$ $10$ $10$ $1$ $1$ $1$
+GroupDRO lr $3\times 10^{-5}$ $3\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
+group lr $0.01$ $0.05$ $0.01$ $1\times 10^{-5}$ $1\times 10^{-5}$ $0.05$ $0.005$
+FedProx lr $3\times 10^{-5}$ $1\times 10^{-3}$ $5\times 10^{-4}$ $1\times 10^{-5}$ $3\times 10^{-5}$ $1\times 10^{-5}$ $8\times 10^{-6}$
+$\mu$ $1\times 10^{-3}$ $0.1$ $1\times 10^{-3}$ $1\times 10^{-3}$ $0.1$ $0.01$ $1\times 10^{-3}$
+Scaffold lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $3\times 10^{-6}$ $8\times 10^{-6}$
+AFL lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$ $1\times 10^{-5}$ $3\times 10^{-6}$ $8\times 10^{-6}$
+FedDG lr $3\times 10^{-5}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
+$\lambda$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$ $U(0,1)$
+FedADG Classifier lr $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
+Gen lr $7\times 10^{-4}$ $5\times 10^{-4}$ $7\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$
+Disc lr $7\times 10^{-4}$ $5\times 10^{-4}$ $7\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-5}$
+$\alpha$ $0.15$ $0.25$ $0.1$ $0.05$ $0.2$
+FedSR lr $1\times 10^{-5}$ $5\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-4}$ $3\times 10^{-5}$ $5\times 10^{-6}$ $5\times 10^{-6}$
+l2 regularizer $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$ $0.01$
+cmi regularizer $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$ $0.001$
+FedGMA lr $3\times 10^{-5}$ $1\times 10^{-4}$ $5\times 10^{-4}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $5\times 10^{-6}$ $1\times 10^{-5}$
+mask-threshold $0.1$ $0.1$ $0.4$ $0.4$ $0.4$ $0.4$ $0.4$
+Appendix E Additional FL-specific challenges for domain generalization
+As mentioned in subsection 4.3, we also include some deeper exploration over the effect of number of clients and communication frequency, which are unique to the FL regime.
+i) Massive number of clients: In this experiment, we explore the performance of different algorithms when the number of clients $C$ increases on PACS.
 We fix the communication rounds $50$ and the local number of epoch is $1$ (synchronizing the models every epoch).
@@
 FedADG and FedSR are are sensitive to the clients number, and they both fail after $C\geq 20.$ 2) Even in the simplest homogeneous setting $\lambda=1,$ where each local client has i.i.d training data, current domain generalization methods IRM, FISH, Mixup, MMD, Coral, GroupDRO work poorly in the existence of large clients number, this means new methods are needed for DG in FL context when data are stored among massive number of clients.
-Figure 2: PACS: Held-out DG test accuracy versus number of clients.
+Figure 2: Performance based on accuracy versus the number of clients across the PACS, CelebA, and Camelyon17 datasets.
+Figure 3: PACS: Held-out DG test accuracy vs. varying communications (resp. varying echoes ).
+Figure 4: Convergence curve on PACS; total clients and training domains $(C,D)=(100,2);$ increasing domain heterogeneity from left to right: $\lambda=(1,0.1,0).$
+Figure 5: Accuracy versus communication rounds for IWildCam; Total clients number $C=243$; increasing heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
 ii) Communication constraint: To show the effect of communication rounds on convergence, we plot the test accuracy versus communication rounds in Appendix Fig.
@@
 The relationship between DG performance and communications requires further exploration.
-Figure 3: PACS: Held-out DG test accuracy vs. varying communications (resp. varying echoes ).
-Appendix D Supplementary Results
+Appendix F Supplementary Results
 In the main paper, we provide experiments results using held-out validation early stopping.
-Here we report the results using in-domain validation set.
+Here we report the results using in-domain validation set for reference.
 We also report the convergence curve for each methods on each dataset for reference.
 We observe that under most of cases, the held-out validation gives us a better model.
-Thus, we recommend using held-out validation set to perform early stopping.
+Thus, we recommend using held-out validation set to perform early stopping when we can access multiple training domains.
+However, if the training dataset contains only 2-3 domains, we should consider using in-domain validation set.
 More results on PACS and IWildCam dataset.
-Here we both report the convergence curve using the DG accuracy of each method on PACS Fig.
-4 and IWildCam Fig.
-5 and the DG accuracy on PACS and IWildCam Table 6 using the in-domain validation.
-From the PACS convergence curve Fig.
-4, we observe that with lower $\lambda$, the model is harder to converge.
-FedAvg ang FedDG converge faster than all other methods while FedADG and FedSR does not converge.
-From IWildCam convergence curve Fig.
-5, we observe that all methods are struggling to converge.
-This is due to the main challenge first come from the $R_{\text{FL}}$ here.
-We thus observe that Federated methods dealing with heterogeneity clients achieves best performance.
-Especially when $\lambda$ is low.
-It is worth noticing that this does not mean DG in FL is solved, none of the methods even achieve the centralized ERM performance.
-Figure 4: Convergence curve on PACS; total clients and training domains $(C,D)=(100,2);$ increasing domain heterogeneity from left to right: $\lambda=(1,0.1,0).$
-Figure 5: Accuracy versus communication rounds for IWildCam; Total clients number $C=243$; increasing heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
-Table 6: Test accuracy on PACS and IWildCam dataset with in-domain validation where FedAvg-ERM is the simple baseline (B).
+Results on FEMNIST dataset.
+As mentioned in the main paper, we include the FEMNIST dataset here for reference.
+It is evident from Table 11 that $\lambda$ does not significantly influence the final domain generalization (DG) accuracy, and whether using in-domain validation or held-out-domain validation does not impact the final DG accuracy as well.
+This suggests a lack of statistical heterogeneity across different domains.
+Furthermore, we observe that changing $\lambda$ does not significantly affect the convergence.
+Most of the results do not converge to the performance level of the centralized counterpart.
+This discrepancy arises from the challenge posed by the large number of clients, where $R_{\text{FL}}=0.980<1$.
+Figure 6: Accuracy versus communication rounds for FEMNIST; total clients and training domains $(K,M)=(100,2586);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
+Figure 7: Accuracy versus communication rounds for Camelyon17; total clients and training domains $(K,M)=(100,4);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
+Figure 8: Accuracy versus communication rounds for CelebA; total clients and training domains $(K,M)=(100,2);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
+Table 8: Test accuracy on PACS and IWildCam dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
 “-” means the method is not applicable in that context.
 Bold is for best and italics is for second best in each column.
+We report the standard deviation among 3 runs.
 PACS ($D=2$) IWildCam ($D=243$)
-$C=1$ $C=100$ (FL) $C=1$ $C=243$ (FL)
+$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
 (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
 B
-FedAvg-ERM $0.8389$ $0.8766$ $\mathit{0.9144}$ $\mathbf{0.8958}$ $\mathit{0.2727}$ $\mathbf{0.1762}$ $\mathbf{0.1241}$ $0.0707$
-DG Adapted IRM $0.9180$ $0.8156$ $0.5449$ $0.1669$ $0.0718$ $0.0903$
-Fish $0.9030$ $0.8497$ $0.6311$ $0.2103$ $0.0781$ $0.0977$
-Mixup $0.8635$ $0.7653$ $0.5551$ $0.1323$ $0.0264$ $0.0370$
-MMD $\mathit{0.9186}$ $0.8150$ $0.6341$ $0.3134$ $0.0718$ $0.0924$
-DeepCoral $\mathbf{0.9216}$ $0.8150$ $0.5515$ $\mathbf{0.3270}$ $0.0718$ $0.0926$
-GroupDRO $0.9060$ $\mathbf{0.9395}$ $\mathbf{0.9437}$ $0.2531$ $0.0530$ $0.0908$
-FL FedProx - $0.8325$ $0.8407$ $0.8568$ $\mathbf{0.2220}$ $\mathbf{0.1770}$ $\mathbf{0.1535}$
-Scaffold - $0.8215$ $0.8643$ $0.8787$ $\mathit{0.2098}$ $\mathit{0.1532}$ $\mathit{0.1467}$
-FDG FedDG - $\mathit{0.8868}$ $0.8880$ $\mathit{0.8952}$ $0.1445$ $0.1023$ $0.0503$
-FedADG - $0.0915$ $0.0732$ $0.0592$ $0.0049$ $0.0049$ $0.0049$
-FedSR - $0.1180$ $0.1341$ $0.1138$ $0.0056$ $0.0056$ $0.0056$
-FedGMA - $0.8820$ $0.8467$ $0.8180$ $0.0271$ $0.0259$ $0.0081$
-More results on CivilComments and Py150 dataset.
-Here we report the DG accuracy on CivilComments and Py150 Table 6 using the in-domain validation.
-We could observe that most of the methods perform well on CivilComments and Py150.
-This could be attributed to the use of pretrained models.
-Additionally, the utilization of pretrained models may explain the relatively high value of $R_{\text{DG}}$, as NLP pretrained models have already been exposed to various domains.
-Table 7: Test accuracy on CivilComments and Py150 datasets with in-domain validation where FedAvg-ERM is the simple baseline (B).
+FedAvg-ERM $\bf 0.943\pm 0.012$ $\bf 0.948\pm 0.020$ $\mathbf{0.955\pm 0.016}$ $\mathbf{0.954\pm 0.018}$ $\mathit{0.343\pm 0.003}$ $\mathbf{0.298\pm 0.002}$ $0.245\pm 0.002$ $\mathbf{0.196\pm 0.010}$
+DG Adapted IRM $0.918\pm 0.007$ $0.920\pm 0.041$ $0.856\pm 0.044$ $0.321\pm 0.004$ $0.196\pm 0.006$ $0.179\pm 0.005$
+Fish $\mathit{0.936\pm 0.021}$ $0.653\pm 0.127$ $0.345\pm 0.107$ $\mathbf{0.354\pm 0.001}$ $0.183\pm 0.004$ $0.154\pm 0.005$
+Mixup $0.918\pm 0.006$ $0.886\pm 0.052$ $0.826\pm 0.030$ $0.331\pm 0.007$ $0.086\pm 0.010$ $0.072\pm 0.007$
+MMD $0.928\pm 0.022$ $0.921\pm 0.041$ $0.855\pm 0.044$ $0.324\pm 0.002$ $0.185\pm 0.005$ $0.163\pm 0.007$
+DeepCoral $0.928\pm 0.007$ $0.920\pm 0.041$ $0.856\pm 0.043$ $0.329\pm 0.008$ $0.185\pm 0.005$ $0.163\pm 0.007$
+GroupDRO $0.928\pm 0.007$ $\mathit{0.944\pm 0.025}$ $0.945\pm 0.015$ $0.211\pm 0.002$ $0.134\pm 0.005$ $0.198\pm 0.006$
+FL FedProx - $0.896\pm 0.025$ $0.891\pm 0.025$ $0.896\pm 0.024$ $0.251\pm 0.003$ $0.204\pm 0.001$ $0.167\pm 0.003$
+Scaffold - $0.896\pm 0.025$ $0.892\pm 0.024$ $0.896\pm 0.024$ $\mathit{0.281\pm 0.002}$ $\mathbf{0.255\pm 0.004}$ $\mathit{0.178\pm 0.008}$
+AFL - $0.931\pm 0.026$ $0.915\pm 0.040$ $0.911\pm 0.041$ $0.256\pm 0.010$ $0.162\pm 0.008$ $0.034\pm 0.001$
+FDG FedDG - $0.933\pm 0.020$ $\mathit{0.947\pm 0.018}$ $\mathit{0.943\pm 0.023}$ $0.274\pm 0.001$ $0.235\pm 0.002$ $0.167\pm 0.004$
+FedADG - $0.943\pm 0.011$ $0.942\pm 0.001$ $0.935\pm 0.011$ $0.259\pm 0.012$ $\mathit{0.251\pm 0.010}$ $0.164\pm 0.001$
+FedSR - $0.640\pm 0.158$ $0.548\pm 0.086$ $0.537\pm 0.092$ $0.177\pm 0.005$ $0.141\pm 0.005$ $0.086\pm 0.003$
+FedGMA - $0.750\pm 0.050$ $0.730\pm 0.087$ $0.724\pm 0.099$ $0.223\pm 0.001$ $0.151\pm 0.002$ $0.091\pm 0.003$
+Table 9: Test accuracy on CelebA and Camelyon17 dataset with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
+“-” means the method is not applicable in that context.
+Bold is for best and italics is for second best in each column.
+We report the standard deviation among 3 runs.
+CelebA ($D=2$) Camelyon17 ($D=3$)
+$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
+(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
+B
+FedAvg-ERM $0.769\pm 0.035$ $0.606\pm 0.019$ $0.517\pm 0.039$ $0.446\pm 0.018$ $0.903\pm 0.009$ $\mathit{0.949\pm 0.007}$ $\mathit{0.950\pm 0.004}$ $\bf 0.948\pm 0.004$
+DG Adapted IRM $\mathbf{0.891\pm 0.012}$ $0.706\pm 0.048$ $0.737\pm 0.013$ $0.912\pm 0.056$ $0.948\pm 0.001$ $0.952\pm 0.002$
+Fish $\mathit{0.883\pm 0.010}$ $0.144\pm 0.158$ $0.072\pm 0.054$ $0.922\pm 0.006$ $0.881\pm 0.006$ $0.878\pm 0.008$
+Mixup $0.288\pm 0.498$ $0.129\pm 0.217$ $0.126\pm 0.218$ $\mathit{0.940\pm 0.011}$ $\mathbf{0.949\pm 0.003}$ $0.947\pm 0.006$
+MMD $0.835\pm 0.047$ $0.809\pm 0.027$ $0.766\pm 0.030$ $0.922\pm 0.027$ $0.946\pm 0.004$ $0.947\pm 0.006$
+DeepCoral $0.852\pm 0.036$ $\mathbf{0.813\pm 0.017}$ $\mathbf{0.782\pm 0.022}$ $\mathbf{0.940\pm 0.007}$ $0.948\pm 0.004$ $0.949\pm 0.003$
+GroupDRO $0.869\pm 0.034$ $\mathit{0.811\pm 0.028}$ $0.761\pm 0.056$ $0.919\pm 0.027$ $0.947\pm 0.006$ $\mathbf{0.953\pm 0.001}$
+FL FedProx - $0.126\pm 0.237$ $0.121\pm 0.245$ $000.000\pm 0$ $0$ $0.939\pm 0.002$ $0.935\pm 0.001$ $\mathit{0.937\pm 0.007}$
+Scaffold - $0.728\pm 0.010$ $\mathit{0.776\pm 0.016}$ $\bf 0.800\pm 0.017$ $0.942\pm 0.001$ $0.929\pm 0.003$ $0.932\pm 0.001$
+AFL - $0.811\pm 0.024$ $0.834\pm 0.005$ $0.828\pm 0.005$ $0.942\pm 0.011$ $0.947\pm 0.006$ $0.932\pm 0.008$
+FDG FedDG - $0.561\pm 0.184$ $0.531\pm 0.112$ $0.464\pm 0.201$ $0.863\pm 0.012$ $0.856\pm 0.003$ $0.869\pm 0.016$
+FedADG - $0.674\pm 0.003$ $0.669\pm 0.008$ $\mathit{0.600\pm 0.009}$ $0.936\pm 0.001$ $0.933\pm 0.004$ $0.926\pm 0.002$
+FedSR - $0.381\pm 0.018$ $0.363\pm 0.041$ $0.383\pm 0.015$ $0.934\pm 0.010$ $0.917\pm 0.009$ $0.925\pm 0.003$
+FedGMA - $0.620\pm 0.012$ $0.615\pm 0.018$ $0.487\pm 0.008$ $0.901\pm 0.006$ $0.854\pm 0.006$ $0.815\pm 0.025$
+Table 10: Test accuracy on CivilComments and Py150 datasets with held-out-domain validation where FedAvg-ERM is the simple baseline (B).
 “-” means the method is not applicable in that context.
@@
 FedDG and FedADG are designed for image dataset and thus not applicable for CivilComments and Py150.
+We report the standard deviation among 3 runs.
 CivilComments ($D=16$) Py150 ($D=5477$)
@@
 B
-FedAvg-ERM $0.5413$ $0.3929$ $0.3390$ $0.3331$ $\mathbf{0.6833}$ $\mathbf{0.6827}$ $0.6501$ $0.6405$
-DG Adapted IRM $\mathit{0.6213}$ $\mathit{0.6153}$ $\mathbf{0.6093}$ $\mathit{0.6771}$ $0.6722$ $0.6379$ $0.6267$
-Fish $0.6038$ $\mathbf{0.6435}$ $\mathit{0.6086}$ $0.6634$ $0.6731$ $0.6464$ $0.6313$
-MMD $0.5981$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6731$ $0.6577$ $\mathit{0.6565}$
-DeepCoral $0.5827$ $0.6121$ $\mathbf{0.6093}$ $0.6558$ $0.6723$ $0.6384$ $0.6236$
-GroupDRO $\mathbf{0.6383}$ $0.6028$ $0.4804$ $0.5127$ $\mathit{0.6824}$ $0.6421$ $0.6391$
-FL FedProx - $0.4762$ $0.3635$ $\mathit{0.3409}$ $0.6730$ $\mathbf{0.6614}$ $\mathbf{0.6573}$
-Scaffold - $0.4932$ $0.4930$ $\mathbf{0.4208}$ $0.6781$ $\mathit{0.6579}$ $0.6364$
-FDG FedSR - $0.3603$ $0.3389$ $0.3188$ $0.0915$ $0.0860$ $0.0587$
-FedGMA - $0.3341$ $0.3157$ $0.3195$ $0.6628$ $0.3737$ $0.3576$
-Results on FEMNIST dataset.
-As mentioned in the main paper, we include the FEMNIST dataset here as reference.
-We could observe from Table 8 that $\lambda$ does not influence the final DG accuracy, and either in-domain validation or held-out-domain validation does not affect the final DG accuracy.
-This indicates the lacking of the statistical heterogeneity across different domains.
-we observe that changing of $\lambda$ does not significantly affect the convergence.
-Most of which does not converge to the centralized counterpart’s performance, this is due to the challenge coming from large number of clients where $R_{\text{FL}}=0.980<1$.
-Figure 6: Accuracy versus communication rounds for FEMNIST; total clients and training domains $(K,M)=(100,2586);$ increasing domain heterogeneity from left to right panel: $\lambda=(1,0.1,0).$
-Table 8: Test accuracy on FEMNIST dataset with in-domain validation and held-out validation where FedAvg-ERM is the simple baseline (B).
+FedAvg-ERM $0.541\pm 0.003$ $0.359\pm 0.028$ $0.347\pm 0.018$ $0.325\pm 0.009$ $\mathbf{0.683}\pm 0.003$ $\mathbf{0.683}\pm 0.000$ $\mathit{0.650\pm 0.000}$ $\mathit{0.641}\pm 0.000$
+DG Adapted IRM $0.641\pm 0.001$ $\mathit{0.587\pm 0.018}$ $0.534\pm 0.038$ $\mathit{0.677}\pm 0.000$ $\mathit{0.655\pm 0.001}$ $\mathbf{0.653\pm 0.000}$ $0.627\pm 0.001$
+Fish $\mathbf{0.671}\pm 0.000$ $0.424\pm 0.158$ $0.340\pm 0.173$ $0.663\pm 0.000$ $0.650\pm 0.000$ $0.648\pm 0.001$ $\mathbf{0.642\pm 0.001}$
+MMD $\mathit{0.652}\pm 0.000$ $\mathbf{0.634\pm 0.012}$ $\mathbf{0.613\pm 0.014}$ $0.656\pm 0.001$ $0.627\pm 0.000$ $0.629\pm 0.000$ $0.610\pm 0.000$
+DeepCoral $0.585\pm 0.000$ $0.515\pm 0.061$ $0.463\pm 0.080$ $0.656\pm 0.003$ $0.651\pm 0.000$ $0.650\pm 0.001$ $0.639\pm 0.000$
+GroupDRO $0.638\pm 0.003$ $0.475\pm 0.014$ $\mathit{0.474\pm 0.003}$ $0.513\pm 0.001$ $0.589\pm 0.000$ $0.603\pm 0.003$ $0.607\pm 0.002$
+FL FedProx - $0.181\pm 0.032$ $0.173\pm 0.033$ $0.170\pm 0.036$ $0.638\pm 0.000$ $0.633\pm 0.000$ $0.612\pm 0.000$
+Scaffold - $0.389\pm 0.015$ $0.376\pm 0.018$ $\mathit{0.332\pm 0.014}$ $0.642\pm 0.000$ $0.638\pm 0.000$ $0.617\pm 0.000$
+AFL - $0.552\pm 0.021$ $0.474\pm 0.014$ $\mathbf{0.435\pm 0.019}$ $0.486\pm 0.000$ $0.485\pm 0.003$ $0.467\pm 0.001$
+FDG FedSR - $0.360\pm 0.003$ $0.339\pm 0.003$ $0.319\pm 0.003$ $0.533\pm 0.002$ $0.526\pm 0.001$ $0.445\pm 0.003$
+FedGMA - $0.209\pm 0.028$ $0.202\pm 0.024$ $0.200\pm 0.022$ $0.620\pm 0.000$ $0.613\pm 0.000$ $0.600\pm 0.000$
+Table 11: Test accuracy on FEMNIST dataset with held-out validation where FedAvg-ERM is the simple baseline (B).
 “-” means the method is not applicable in that context.
 Bold is for best and italics is for second best in each column.
-FEMNIST (In-Domain) FEMNIST (Held-Out-Domain)
-$C=1$ $C=100$ (FL) $C=1$ $C=100$ (FL)
-(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$ (centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
+FEMNIST (Held-Out-Domain)
+$C=1$ $C=100$ (FL)
+(centralized) $\lambda=1$ $\lambda=0.1$ $\lambda=0$
 B
-FedAvg-ERM $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$ $\mathbf{0.8544}$ $0.8368$ $0.8398$ $0.8399$
-DG Adapted IRM $0.8442$ $0.8303$ $0.8389$ $0.8328$ $0.8442$ $0.8303$ $0.8389$ $0.8328$
-Fish $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$ $\mathit{0.8487}$ $0.8331$ $0.8290$ $0.8264$
-Mixup $0.8337$ $0.8275$ $0.8168$ $0.8161$ $0.8337$ $0.8275$ $0.8168$ $0.8161$
-MMD $0.8442$ $0.8430$ $0.8323$ $0.8293$ $0.8442$ $0.8430$ $0.8323$ $0.8293$
-DeepCoral $0.8457$ $0.8403$ $0.8328$ $0.8513$ $0.8457$ $0.8403$ $0.8328$ $0.8513$
-GroupDRO $0.8413$ $0.8447$ $0.8139$ $0.8052$ $0.8413$ $0.8447$ $0.8139$ $0.8052$
-FL FedProx - $0.8448$ $0.8440$ $0.8425$ $0.8448$ $0.8440$ $0.8425$
-Scaffold - $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$ $0.8531$ $\mathit{0.8507}$ $\mathit{0.8575}$
-FDG FedDG - $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$ $\mathbf{0.8591}$ $\mathbf{0.8573}$ $\mathbf{0.8576}$
-FedADG - $0.0549$ $0.0557$ $0.0548$ $0.0549$ $0.0557$ $0.0548$
-FedSR - $0.8320$ $0.8316$ $0.8306$ $0.8320$ $0.8316$ $0.8306$
-FedGMA - $\mathit{0.8486}$ $0.8395$ $0.8371$ $0.8486$ $0.8395$ $0.8371$
-Appendix E Gap Table
+FedAvg-ERM $\mathbf{0.854\pm 0.001}$ $0.837\pm 0.003$ $0.838\pm 0.002$ $0.834\pm 0.001$
+DG Adapted IRM $0.844\pm 0.001$ $0.832\pm 0.007$ $0.822\pm 0.005$ $0.821\pm 0.004$
+Fish $\mathit{0.849\pm 0.002}$ $0.833\pm 0.004$ $0.829\pm 0.003$ $0.826\pm 0.008$
+Mixup $0.834\pm 0.002$ $0.828\pm 0.009$ $0.821\pm 0.008$ $0.816\pm 0.005$
+MMD $0.844\pm 0.003$ $0.843\pm 0.006$ $0.832\pm 0.007$ $0.829\pm 0.006$
+DeepCoral $0.846\pm 0.002$ $0.840\pm 0.005$ $0.836\pm 0.002$ $0.851\pm 0.003$
+GroupDRO $0.841\pm 0.004$ $0.835\pm 0.005$ $0.814\pm 0.002$ $0.805\pm 0.003$
+FL FedProx - $\mathbf{0.865\pm 0.001}$ $\mathbf{0.864\pm 0.001}$ $\mathbf{0.863\pm 0.001}$
+Scaffold - $\mathit{0.860\pm 0.002}$ $\mathit{0.859\pm 0.002}$ $\mathit{0.858\pm 0.002}$
+AFL - $0.629\pm 0.010$ $0.593\pm 0.037$ $0.584\pm 0.020$
+FDG FedDG - $0.859\pm 0.002$ $0.857\pm 0.002$ $0.856\pm 0.002$
+FedADG - $0.817\pm 0.026$ $0.800\pm 0.061$ $0.785\pm 0.025$
+FedSR - $0.832\pm 0.015$ $0.832\pm 0.020$ $0.831\pm 0.014$
+FedGMA - $0.849\pm 0.010$ $0.842\pm 0.004$ $0.837\pm 0.003$
+Appendix G Gap Table
 We list the gap table in Table 12 for summarizing the current DG algorithms performance gap w.r.t FedAvg-ERM in the FL context, in particular, positive means it outperforms FedAvg-ERM, negative means it is worse than FedAvg-ERM.
 It can be seen that in the on the simple dataset, the best DG migrated from centralized setting is better than FedAvg-ERM.
-In the domain separation case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting.
+In the complete heterogeneity case, no centralized DG algorithms can be adapted to it, and FDG methods performs comparably good in this setting.
 However, they fail in harder datasets.
@@
 This highlights the need for future research and development of DG methods in the FL regime.
-Table 9: Gap Table: Current progress in solving Federated DG with respect to FedAvg-ERM.
-Domain Separation
+Table 12: Gap Table: Current Progress in solving DG in FL context
+Domain Seperation
 Free of Data Leakage
@@
 IWildCam
+CelebA
+Camelyon17
 CivilComments
 Py150
-Adapted centralized methods $+0.0629$ $-0.0981$ $+0.0421$ $-0.0062$
-Federated methods $+0.0060$ $+0.0458$ $+0.0421$ $-0.0062$
-FDG FedDG $+0.0468$ $-0.0775$
-FedADG $-0.6874$ $-0.2171$
-FedSR $-0.7520$ $-0.2164$ $-0.0326$ $-0.5912$
-FedGMA $+0.0054$ $-0.1263$ $-0.0588$ $-0.0199$
-Appendix F Training Time, Communication Rounds and local computation
+Centralized adopted methods $-0.010$ $-0.047$ $+0.265$ $+0.003$ $+0.266$ $+0.003$
+Federated methods $-0.040$ $+0.010$ $+0.319$ $-0.003$ $+0.266$ $-0.012$
+FDG FedDG $-0.008$ $-0.010$ $+0.014$ $-0.094$
+FedADG $-0.013$ $+0.060$ $+0.152$ $-0.017$
+FedSR $-0.407$ $-0.104$ $-0.154$ $-0.033$ $-0.008$ $-0.124$
+FedGMA $-0.225$ $-0.094$ $+0.098$ $-0.096$ $-0.145$ $-0.037$
+Appendix H Training Time, Communication Rounds and local computation
 In this section, we provide training time per communication in terms of the wall clock training time.
 Notice that for a fixed dataset, most of algorithms have similar training time comparing to FedAvg-ERM, where FedDG and FedADG are significantly more expensive.
-Table 10: Wall-clock Training time per communication (unit: s).
-PACS IWildCam CivilComments Py150 FEMNIST
-$C=100$ $C=243$ $C=100$ $C=100$ $C=100$
+Table 13: Wall-clock Training time per communication (unit: s). There might be variance due to the machine’s status and resource availability.
+PACS IWildCam CelebA Camelyon17 CivilComments Py150 FEMNIST
+$C=100$ $C=243$ $C=100$ $C=100$ $C=100$ $C=100$ $C=100$
 B
-FedAvg-ERM $143$ $6301$ $3958$ $6566$ $262$
-DG Adapted IRM $147$ $6454$ $4085$ $7089$ $297$
-Fish $148$ $7072$ $5483$ $7770$ $324$
-Mixup $144$ $6294$ $264$
-MMD $144$ $6663$ $4024$ $7603$ $287$
-DeepCoral $144$ $6597$ $3901$ $7212$ $287$
-GroupDRO $145$ $9311$ $4690$ $8121$ $307$
-FL FedProx $169$ $6921$ $4502$ $6513$ $288$
-Scaffold $167$ $6876$ $4421$ $6389$ $281$
-FDG FedDG $352$ $32172$ $989$
-FedADG $181$ $11094$ $907$
-FedSR $151$ $7136$ $4403$ $8020$ $280$
-FedGMA $143$ $6795$ $4525$ $6545$ $261$
+FedAvg-ERM $143$ $6301$ $298$ $845$ $3958$ $6566$ $262$
+DG Adapted IRM $147$ $6454$ $309$ $1163$ $4085$ $7089$ $297$
+Fish $148$ $7072$ $312$ $1003$ $5483$ $7770$ $324$
+Mixup $144$ $6294$ $302$ $933$ $264$
+MMD $144$ $6663$ $312$ $963$ $4024$ $7603$ $287$
+DeepCoral $144$ $6597$ $313$ $879$ $3901$ $7212$ $287$
+GroupDRO $145$ $9311$ $310$ $750$ $4690$ $8121$ $307$
+FL FedProx $169$ $6921$ $1219$ $4310$ $4502$ $6513$ $288$
+Scaffold $167$ $6876$ $1344$ $4623$ $4421$ $6389$ $281$
+FDG FedDG $352$ $32172$ $9232$ $15784$ $989$
+FedADG $181$ $11094$ $6502$ $9945$ $907$
+FedSR $151$ $7136$ $2267$ $6567$ $4403$ $8020$ $280$
+FedGMA $143$ $6795$ $558$ $2036$ $4525$ $6545$ $261$